### v1

# Super-extensions of tensor algebras and their applications

Abstract.
Following [It12] and [It15], we construct two super-extensions of the usual tensor algebra through the super-actions of symmetric groups and Hecke algebras respectively. For each extension, we consider a special type of derivations coming from covectors, and study the the space generated, in some special manner, by these derivations and operators from left multiplication by vectors and permutations. Duality theorems of these spaces and the super-actions are proved. As an application, we provide a new proof of the Schur-Sergeev duality theorem, as well as its quantum version.
1. Introduction
Denote by $V^{\otimes d}$ the $d$-fold of a vector space $V$. In [Sc27], Schur considered the relation between the permutation action of the symmetric group $\mathfrak{S}_{d}$ and the diagonal action of the general linear group $\mathrm{GL(V)}$ on $V^{\otimes d}$. This is the well-known Schur-Weyl duality nowadays. It plays a fundamental role in representation theory (see, e.g., [We46]). Since its importance, this duality theorem has been generalized in several directions during the past decades. For example, in the late 1980s, Sergeev [Ser85] and Berele and Regev [BR87] discovered independently a super version of the Schur-Weyl duality. Meanwhile, Jimbo [Jim86] introduced a duality theorem about Hecke algebras and quantum enveloping algebras. In [Mi06], Mitsuhashi extended Jimbo’s theorem to the super case.
Usually, in order to prove these dualities, people use the double commutant theorem. For example, for the classical Schur-Weyl duality, one needs to show that $\mathrm{GL}_{d}(\mathbb{C})=\mathrm{End}_{\mathfrak{S}_{d}}(\mathbb{C}^{d})$. Then by the semisimplicity of the group algebra $\mathbb{C}\mathfrak{S}_{d}$ one has $\mathbb{C}\mathfrak{S}_{d}=\mathrm{End}_{\mathrm{GL}_{d}(\mathbb{C})}(\mathbb{C}^{d})$. The verification of the inclusion $\mathrm{GL}_{d}(\mathbb{C})\subset\mathrm{End}_{\mathfrak{S}_{d}}(\mathbb{C}^{d})$ is trivial. But the reverse direction requires further non-trivial techniques. We refer the reader to [GW03] and [CW12] for detailed proof.
In [It12], Itoh generalized the usual tensor algebra $T(V)$ by considering actions of symmetric groups. He studied operators of left multiplication by vectors and linear functionals, and established a duality theorem between the space generated by these operators and the infinity symmetric group. As a consequence, he provided a new proof of the Schur-Weyl dual theorem. In a subsequent paper [It15], he generalized his constructions to quantum case in which symmetric group actions are replaced by Hecke algebra actions. Using a similar argument, he gave a new proof of Jimbo’s duality theorem.
The main purpose of this paper is to generalize Itoh’s results to super cases. First of all, we consider the right super-action of $\mathfrak{S}_{d}$ on $d$-fold of a vector superspace $V=V_{\overline{0}}\oplus V_{\overline{1}}$. Using this action, we define an associative algebra structure on the space
$\overline{T}(V)=\bigoplus_{d\geq 0}V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathfrak{S}_{\infty}.$
For each covector $v^{\ast}\in V^{\ast}$, we define a derivation $L(v^{\ast})$ of $\overline{T}(V)$. Denote by $L(w)$ and $L(\sigma)$ the operators coming from left multiplication by a vector $w$ and a permutation $\sigma$ respectively. We show that the space $\mathcal{A}_{d}^{k}\subset\mathrm{End}_{\mathbb{C}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$ spanned by operators of the form
$L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{\ast})\cdots L(v_{d}^{\ast}),$
is the commutant of $\mathfrak{S}_{d+k}$. Notice that $\mathcal{A}_{d}^{0}$ is just action of the Lie superalgebra $\mathfrak{gl}(m|n)$. We provide a new proof of the Schur-Sergeev duality. Secondly, we consider the right super-action of the Hecke algebra $\mathscr{H}_{d}$ on $V^{\otimes d}$, and construct an algebra
$\widetilde{T}(V)=\bigoplus_{d\geq 0}V^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{\infty}.$
Similar duality in this algebra is proved. As an application, we provide a new proof of Mitsuhashi’s duality.
This paper is organized as follows. We construct the super-extension $\overline{T}(V)$ in Section 2 and study the communication relations among $L(v^{\ast})$, $L(w)$, and $L(\sigma)$ in Section 3. Duality theorems are proved in Section 4. We deal with the the quantum version $\widetilde{T}(V)$ in Section 5, 6, and 7.
Acknowledgement
We would like to thank Li Luo for many helpful discussions. Especially, he introduced the Schur-Sergeev duality and its quantum version to us which initiate this work.
2. Super-extensions through symmetric groups
2.1. Symmetric groups
For any positive integer $d$, we denote by $\mathfrak{S}_{d}$ the symmetric group of the set $[d]=\{1,2,\ldots,d\}$, and by $s_{i}$, $11\leq i\leq d-$ 1, the transposition interchanging $i$ and $i+1$. We often use cycle notation to denote a permutation. So $\tau=(i_{1}\ i_{2}\ \cdots\ i_{k})\in\mathfrak{S}_{d}$ maps $i_{r}$ to $i_{r+1}$ for $11\leq r\leq k-$ 1 and $i_{r}$ to $i_{1}$, and fixes other numbers. There is a natural sequence of inclusions: $\mathfrak{S}_{1}\subset\cdots\subset\mathfrak{S}_{d}\subset\cdots$. The direct limit is denoted by $\mathfrak{S}_{\infty}$. One can view $\mathfrak{S}_{\infty}$ as the group generated by $\{s_{i}\}_{i\geq 1}$ subject to the relations:
$s_{i}^{2}=1,$
$s_{i}s_{i+1}s_{i}=s_{i+1}s_{i}s_{i+1},$
and
$s_{i}s_{j}=s_{j}s_{i},\text{ whenever }|i-j|>1.$
Given $k\geq 1$, we define a group homomorphism :${}^{\uparrow k}:\mathfrak{S}_{\infty}\rightarrow\mathfrak{S}_{\infty}$ by $s_{i}^{\uparrow k}=s_{i+k}$.
2.2. Superspaces and general linear Lie superalgebras
Let $m,n$ be two fixed positive integers, and $V=\mathbb{C}^{m|n}=\mathbb{C}^{m}\oplus\mathbb{C}^{n}$. As usual, we set $V_{\overline{0}}=\mathbb{C}^{m}$ and $V_{\overline{1}}=\mathbb{C}^{n}$. Vectors in $V_{\overline{0}}$ (resp. $V_{\overline{1}}$) are of parity 00 (resp. parity $1$). The parity of a homogeneous element $v\in V$ is denoted by $|v|$. In this section, we will denote by $e_{i}$, $1\leq i\leq m+n$, the vector in $V$ that is 1 in the $i$-th position and 0 elsewhere. Then $\{e_{1},\ldots,e_{m}\}$ and $\{e_{m+1},\ldots,e_{m+n}\}$ are bases of $V_{\overline{0}}$ and $V_{\overline{1}}$ respectively. We denote by $\{e_{1}^{\ast},\ldots,e_{m+n}^{\ast}\}$ the dual basis of $\{e_{1},\ldots,e_{m+n}\}$, i.e., $e_{i}^{\ast}(e_{j})=\delta_{ij}$. Here $\delta_{ij}$ is the Kronecker symbol. Define a parity function on this dual basis by requiring that $|e_{i}^{\ast}|=|e_{i}|$ for each $1\leq i\leq m+n$. Thus the dual space $V^{\ast}=V_{\overline{0}}^{\ast}\oplus V_{\overline{1}}^{\ast}$ of $V$ is a $\mathbb{Z}_{2}$-graded vector space.
The general linear Lie superalgebra $\mathfrak{gl}(m|n)$ is
$\bigg{\{}\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\bigg{|}A\in M_{m\times m}(\mathbb{C}),B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg{\}}.$
We can see that
$\mathfrak{gl}(m|n)=\mathfrak{gl}(m|n)_{\overline{0}}\oplus\mathfrak{gl}(m|n)_{\overline{1}},$
where
$\mathfrak{gl}(m|n)_{\overline{0}}=\bigg{\{}\left(\begin{array}[]{cc}A&0\\ 0&D\end{array}\right)\bigg{|}A\in M_{m\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg{\}},$
and
$\mathfrak{gl}(m|n)_{\overline{1}}=\bigg{\{}\left(\begin{array}[]{cc}0&B\\ C&0\end{array}\right)\bigg{|}B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C})\bigg{\}}.$
The parity of a homogeneous element $g\in\mathfrak{gl}(m|n)$ is denoted by $|g|$. As usual, for any $1\leq i,j\leq m+n$, let $E_{ij}$ be the $(m+n)\times(m+n)$ matrix with $1$ in the $(i,j)$-position and 00 elsewhere. Then $|E_{ij}|=|e_{i}|+|e_{j}^{\ast}|$.
2.3. Representations on $V^{\otimes d}$
For further discussion, we make some conventions. From now on, elements in $V$ or $\mathfrak{gl}(m|n)$ appeared in our paper are always assumed to be homogeneous unless otherwise specified. Subscripts are often arranged in a decreasing order. For example, a pure $d$-tensor of $V$ is written as $v_{d}\otimes\cdots\otimes v_{1}$.
The general linear Lie superalgebra $\mathfrak{gl}(m|n)$ acts from left on $V$ in a natural way. This action induces a representation $\rho_{d}$ of the enveloping algebra $U(\mathfrak{gl}(m|n))$ on $V^{\otimes d}$ for each $d\geq 1$ as follows:
$\rho_{d}(g)(v_{d}\otimes\cdots\otimes v_{1})=\sum_{k=1}^{d}(-1)^{|g|(|v_{d}|+\cdots+|v_{k+1}|)}v_{d}\otimes\cdots\otimes gv_{k}\otimes\cdots\otimes v_{1},$
where $g\in\mathfrak{gl}(m|n)$ and $v_{1},\ldots,v_{d}\in V$.
On the other hand, the symmetric group $\mathfrak{S}_{d}$ acts from right on $[m+n]^{d}=\{(i_{d},\ldots,i_{1})|1\leq i_{1},\ldots,i_{d}\leq m+n\}$ as follows: for each $1\leq k<d$ and $I\in[m+n]^{d}$,
$I.s_{k}=(i_{d},\ldots,i_{k},i_{k+1},\ldots,i_{1}).$
Thus $I.\sigma=(i_{\sigma(d)},\ldots,i_{\sigma(1)})$ for $\sigma\in\mathfrak{S}_{d}$. We then have a right $\mathfrak{S}_{d}$-action on $V^{\otimes d}$ defined by
$(v_{d}\otimes\cdots\otimes v_{1}).s_{i}=(-1)^{|v_{i}|\cdot|v_{i+1}|}v_{d}\otimes\cdots\otimes v_{i}\otimes v_{i+1}\otimes\cdots\otimes v_{1}.$
For example, if $\sigma=(123)=s_{1}s_{2}$, then
$(v_{3}\otimes v_{2}\otimes v_{1}).\sigma=(-1)^{|v_{2}|\cdot|v_{1}|+|v_{3}|\cdot|v_{1}|}v_{1}\otimes v_{3}\otimes v_{2}.$
In general, for any $\sigma\in\mathfrak{S}_{d}$,
(2.1) $(v_{d}\otimes v_{d-1}\otimes\cdots\otimes v_{1}).\sigma=\varepsilon_{v_{d},\ldots,v_{1};\sigma}v_{\sigma(d)}\otimes v_{\sigma(d-1)}\otimes\cdots\otimes v_{\sigma(1)},$
where $\varepsilon_{v_{d},\ldots,v_{1};\sigma}$ is either $1$ or $-1$, depending on $v_{d},\ldots,v_{1}$ and $\sigma$.
2.4. Super-extensions
From now on, we will write a pure tensor $v_{d}\otimes\cdots\otimes v_{1}\in V^{\otimes d}$ shortly as $v_{d}\cdots v_{1}$. We denote
$\overline{T}_{d}(V)=V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{\infty},$
and
$\overline{T}(V)=\bigoplus_{d\geq 0}\overline{T}_{d}(V).$
We define an associative product $\cdot$ on $\overline{T}$ as follows: for any $u_{k}\cdots u_{1}\otimes\tau\in\widehat{T}_{k}(V)$ and $v_{d}\cdots v_{1}\otimes\sigma\in\overline{T}_{d}(V)$,
$(u_{k}\cdots u_{1}\otimes\tau)\cdot(v_{d}\cdots v_{1}\otimes\sigma)=u_{k}\cdots u_{1}v_{d}\cdots\otimes v_{1}\otimes\tau^{\uparrow d}\sigma.$
For any $v,w\in V$, we have immediately that
$(vw)\cdot s_{1}=(-1)^{|v|\cdot|w|}wv,$
and
$s_{i}\cdot v=v\otimes s_{i+1}.$
For any $\varphi\in\overline{T}(V)$, we define
$\begin{array}[]{cccc}L(\varphi):&\overline{T}(V)&\rightarrow&\overline{T},\\[3.0pt] &\psi&\mapsto&\varphi\cdot\psi.\end{array}$
For a homogeneous element $v^{\ast}\in V^{\ast}$, we define $L(v^{\ast}):\overline{T}(V)\rightarrow\overline{T}(V)$ by
$L(v^{\ast})(v_{d}\cdots v_{1}\otimes\sigma)$
$=\sum_{k=1}^{d}(-1)^{|v^{\ast}|(|v_{d}|+\cdots+|v_{k+1}|)}\langle v^{\ast},v_{k}\rangle v_{d}\cdots\hat{v}_{k}\cdots v_{1}\otimes(d\ d-1\ \cdots\ k+1\ k)\sigma,$
$=\sum_{k=1}^{d}\Big{(}(v^{\ast}\otimes\mathrm{id}_{V^{\otimes d-1}})\big{(}(v_{d}\cdots v_{1}).(s_{k}\cdots s_{d-2}s_{d-1})\big{)}\Big{)}\otimes s_{d-1}s_{d-2}\cdots s_{k}\sigma$
where the symbol $\hat{v}_{k}$ means the omission of $v_{k}$. The well-definedness of these two maps can be checked directly and stated formally as:
Proposition 2.1.
The operators $L(\sigma)$, $L(e_{i})$, and $L(e_{i}^{\ast})$ for $\sigma\in\mathfrak{S}_{\infty},1\leq i\leq m+n$ are elements in $\mathrm{End}_{\mathfrak{S}_{\infty}}(\overline{T}(V))$, where $\mathfrak{S}_{\infty}$ acts on $\overline{T}(V)$ by right multiplication.
3. Properties of operators
It is easy to verify the following commutation relations.
Proposition 3.1.
For any $1\leq i,j\leq m+n$ and $\sigma\in\mathfrak{S}_{\infty}$, we have
$L(e_{i})L(e_{j})=(-1)^{|e_{i}|\cdot|e_{j}|}L(e_{j})L(e_{i})L(s_{1}),$
$L(e_{i}^{\ast})L(e_{j}^{\ast})=(-1)^{|e_{i}^{\ast}|\cdot|e_{j}^{\ast}|}L(s_{1})L(e_{j}^{\ast})L(e_{i}^{\ast}),$
$L(e_{j}^{\ast})L(e_{i})=(-1)^{|e_{j}|\cdot|e_{i}|}L(e_{i})L(s_{1})L(e_{j}^{\ast})+\langle e_{j}^{\ast},e_{i}\rangle,$
$L(\sigma)L(e_{i})=L(e_{i})L(\sigma^{\uparrow 1}),$
and
$L(e_{i}^{\ast})L(\sigma)=L(\sigma^{\uparrow 1})L(e_{i}^{\ast}).$
The action of $U(\mathfrak{gl}(m|n))$ on $V^{\otimes d}$ can be naturally extended to the action on $\overline{T}_{d}(V)$, and it can be expressed in terms of $L(e_{i})$ and $L(e_{i}^{\ast})$. By direct calculation, we have
$L(e_{i})L(e_{j}^{\ast})(v_{d}\otimes\cdots\otimes v_{1}\otimes\sigma)$
$=\sum_{k=1}^{d}(-1)^{|e_{j}^{\ast}|(|v_{d}|+\cdots+|v_{k+1}|)}\langle e_{j}^{\ast},v_{k}\rangle e_{i}v_{d}\cdots\hat{v}_{k}\cdots v_{1}\otimes s_{d-1}s_{d-2}\cdots s_{k}\sigma$
$=\sum_{k=1}^{d}(-1)^{(|e_{i}|+|e_{j}^{\ast}|)(|v_{d}|+\cdots+|v_{k+1}|)}\langle e_{j}^{\ast},v_{k}\rangle v_{d}\cdots v_{k+1}e_{i}v_{k-1}\cdots v_{1}\otimes\sigma$
$=\sum_{k=1}^{d}(-1)^{(|e_{i}|+|e_{j}^{\ast}|)(|v_{d}|+\cdots+|v_{k+1}|)}v_{d}\cdots v_{k+1}\langle e_{j}^{\ast},v_{k}\rangle e_{i}v_{k-1}\cdots v_{1}\otimes\sigma$
$=\sum_{k=1}^{d}(-1)^{|E_{ij}|(|v_{d}|+\cdots+|v_{k+1}|)}v_{d}\cdots v_{k+1}(E_{ij}v_{k})v_{k-1}\cdots v_{1}\otimes\sigma$
$=\big{(}\rho_{d}(E_{ij})(v_{d}\cdots v_{1})\big{)}\otimes\sigma.$
In particular, on $V^{\otimes d}$, we have
(3.1) $\rho_{d}(E_{ij})=L(e_{i})L(e_{j}^{\ast}).$
The stabilizer of $I\in[m+n]^{d}$ is denoted by $(\mathfrak{S}_{d})_{I}$, i.e., $(\mathfrak{S}_{d})_{I}=\{\sigma\in\mathfrak{S}_{d}|I.\sigma=I\}$. For $1\leq k\leq m+n$, denote
$\mathbf{m}_{k}(I)=\sharp\{j\in[d]~{}|~{}i_{j}=k\},$
$\mathbf{m}(I)=(\mathbf{m}_{1}(I),\ldots,\mathbf{m}_{m+n}(I)),$
and
$\mathbf{m}(I)!=\mathbf{m}_{1}(I)!\cdots\mathbf{m}_{m+n}(I)!.$
For example, $\mathbf{m}_{1}(3,2,3)=0$, $\mathbf{m}_{3}(3,2,3)=2$, and $\mathbf{m}(3,2,3)=(0,1,2,0,\ldots,0)$. Obviously, one has $|(\mathfrak{S}_{d})_{I}|=\mathbf{m}(I)!$.
Lemma 3.2.
Set
$\mathcal{S}_{1}$ =$=\{1,s_{1}\},$
$\mathcal{S}_{2}$ =$=\{1,s_{2},s_{2}s_{1}\},$
$\mathcal{S}_{3}$ =$=\{1,s_{3},s_{3}s_{2},s_{3}s_{2}s_{1}\}$
$\vdots$
$\mathcal{S}_{d-1}$ =$=\{1,s_{d-1},s_{d-1}s_{d-2},\ldots,s_{d-1}s_{d-2}\cdots s_{2}s_{1}\}.$
Then every $\sigma\in\mathfrak{S}_{d}$ can be written uniquely as
$\sigma=\sigma_{1}\sigma_{2}\cdots\sigma_{d-1},$
with $\sigma_{i}\in\mathcal{S}_{i}$. Moreover, the above expression is reduced.
Proof.
For a proof, one can consult [KT08, Corollary 4.4].∎
Proposition 3.3.
For $I=(i_{d},\ldots,i_{1})\in[m+n]^{d}$ and $\tau\in\mathfrak{S}_{d}$, we have
$L(e_{i_{\tau(d)}})\cdots L(e_{i_{\tau(1)}})L(e_{i_{\tau(1)}}^{\ast})\cdots L(e_{i_{\tau(d)}}^{\ast})(e_{i_{d}}\cdots e_{i_{1}})=\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}}.$
Proof.
We have
$L(e_{i_{\tau(1)}}^{\ast})\cdots L(e_{i_{\tau(d)}}^{\ast})(e_{i_{d}}\cdots e_{i_{1}})$
$=L(e_{i_{\tau(1)}}^{\ast})\cdots L(e_{i_{\tau(d-1)}}^{\ast})(e_{i_{\tau(d)}}^{\ast}\otimes\mathrm{id}_{V^{\otimes d-1}})\Bigg{(}\sum_{\sigma_{d-1}\in\mathcal{S}_{d-1}}(e_{i_{d}}\cdots e_{i_{1}}).\sigma_{d-1}^{-1}\otimes\sigma_{d-1}\Bigg{)}$
$=\sum_{\sigma_{d-1}\in\mathcal{S}_{d-1}}L(e_{i_{\tau(1)}}^{\ast})\cdots L(e_{i_{\tau(d-2)}}^{\ast})$
$\ \ \ \ \ \ \ \ \ \ \ \ (e_{i_{\tau(d)}}^{\ast}\otimes e_{i_{\tau(d-1)}}^{\ast}\otimes\mathrm{id}_{V^{\otimes d-2}})\Bigg{(}\sum_{\sigma_{2}\in\mathcal{S}_{d-2}}\big{(}(e_{i_{d}}\cdots e_{i_{1}}).\sigma_{d-1}^{-1}\big{)}.\sigma_{d-2}^{-1}\otimes\sigma_{d-2}\sigma_{d-1}\Bigg{)}$
=$\cdots$= ⋯
$=\sum_{\sigma_{1}\in\mathcal{S}_{1}}\cdots\sum_{\sigma_{d-1}\in\mathcal{S}_{d-1}}(e_{i_{\tau(d)}}^{\ast}\cdots e_{i_{\tau(1)}}^{\ast})\big{(}(e_{i_{d}}\cdots e_{i_{1}}).(\sigma_{d-1}^{-1}\cdots\sigma_{1}^{-1})\big{)}\otimes\sigma_{1}\cdots\sigma_{d-1}$
$=\sum_{\sigma\in\mathfrak{S}_{d}}(e_{i_{\tau(d)}}^{\ast}\cdots e_{i_{\tau(1)}}^{\ast})\big{(}(e_{i_{d}}\cdots e_{i_{1}}).\sigma^{-1}\big{)}\otimes\sigma$
$=\sum_{\sigma\in\mathfrak{S}_{d}}\varepsilon_{e_{i_{d}},\ldots,e_{i_{1}};\sigma^{-1}}(e_{i_{\tau(d)}}^{\ast}\cdots e_{i_{\tau(1)}}^{\ast})(e_{i_{\sigma^{-1}(d)}}\cdots e_{i_{\sigma^{-1}(1)}})\otimes\sigma$
$=\sum_{\sigma\in\mathfrak{S}_{d}}\varepsilon_{e_{i_{d}},\ldots,e_{i_{1}};\sigma^{-1}}\delta_{i_{\tau(d)}i_{\sigma^{-1}(d)}}\cdots\delta_{i_{\tau(1)}i_{\sigma^{-1}(1)}}\sigma$
$=\sum_{\sigma\in\mathfrak{S}_{d}}\varepsilon_{e_{i_{d}},\ldots,e_{i_{1}};\sigma^{-1}}\delta_{i_{\tau(d)}i_{(\tau\sigma)^{-1}\tau(d)}}\cdots\delta_{i_{\tau(1)}i_{(\tau\sigma)^{-1}\tau(1)}}\sigma$
$=\sum_{(\tau\sigma)^{-1}\in(\mathfrak{S}_{d})_{I.\tau}}\varepsilon_{e_{i_{d}},\ldots,e_{i_{1}};\sigma^{-1}}\sigma,$
where the fifth equality follows from Lemma 3.2. Therefore
$L(e_{i_{\tau(d)}})\cdots L(e_{i_{\tau(1)}})L(e_{i_{\tau(1)}}^{\ast})\cdots L(e_{i_{\tau(d)}}^{\ast})(e_{i_{d}}\cdots e_{i_{1}})$
$=\sum_{(\tau\sigma)^{-1}\in(\mathfrak{S}_{d})_{I.\tau}}\varepsilon_{e_{i_{d}},\ldots,e_{i_{1}};\sigma^{-1}}(e_{i_{\tau(d)}}\cdots e_{i_{\tau(1)}}).\sigma$
$=\sum_{(\tau\sigma)^{-1}\in(\mathfrak{S}_{d})_{I.\tau}}\varepsilon_{e_{i_{d}},\ldots,e_{i_{1}};\sigma^{-1}}(e_{i_{\sigma^{-1}(d)}}\cdots e_{i_{\sigma^{-1}(1)}}).\sigma$
$=\sum_{(\tau\sigma)^{-1}\in(\mathfrak{S}_{d})_{I.\tau}}((e_{i_{d}}\cdots e_{i_{1}}).\sigma^{-1}).\sigma$
$=|(\mathfrak{S}_{d})_{I.\tau}|e_{i_{d}}\cdots e_{i_{1}}$
$=|(\mathfrak{S}_{d})_{I}|e_{i_{d}}\cdots e_{i_{1}}$
$=\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}}.$
Corollary 3.4.
For any $I=(i_{d},\ldots,i_{1}),J=(j_{d},\ldots,j_{1})\in[m+n]^{d}$, we have
$L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(e_{j_{d}}\cdots e_{j_{1}})=\begin{cases}\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}},&\text{if }\mathbf{m}(I)=\mathbf{m}(J),\\[3.0pt] 0,&\text{if }\mathbf{m}(I)\neq\mathbf{m}(J).\end{cases}$
We provide two more interesting formulas of such type which will be used in the sequel.
Proposition 3.5.
For any $\varphi\in\overline{T}_{d}(V)$, we have
(3.2) $\sum_{i=1}^{m+n}L(e_{i})L(e_{i}^{\ast})(\varphi)=d\varphi.$
Proof.
For any $(j_{d},\ldots,j_{1})\in[m+n]^{d}$ and $\sigma\in\mathfrak{S}_{\infty}$, we have
$\sum_{i=1}^{m+n}L(e_{i})L(e_{i}^{\ast})(e_{j_{d}}\cdots e_{j_{1}}\otimes\sigma)$ $=\sum_{i=1}^{m+n}\big{(}\rho_{d}(E_{ii})(e_{j_{d}}\cdots e_{j_{1}})\big{)}\otimes\sigma$
$=\rho_{d}\Bigg{(}\sum_{i=1}^{m+n}E_{ii}\Bigg{)}(e_{j_{d}}\cdots e_{j_{1}})\otimes\sigma$
$=d(e_{j_{d}}\cdots e_{j_{1}}\otimes\sigma).$
Since each $\varphi\in\overline{T}_{d}(V)$ is a linear combination of $e_{j_{d}}\cdots e_{j_{1}}\otimes\sigma$’s, we get the result.∎
We define the Euler operator
(3.3) $A_{d}=\frac{1}{d!}\sum_{(i_{d},\ldots,i_{1})\in[m+n]^{d}}L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast}).$
Proposition 3.6.
For any $\varphi\in\overline{T}_{d}(V)$, we have
(3.4) $A_{d}(\varphi)=\varphi.$
Proof.
We have
$A_{d}(\varphi)$ $=\frac{1}{d!}\sum_{(i_{d},\ldots,i_{1})\in[m+n]^{d}}L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)$
$=\frac{1}{d!}\sum_{(i_{d},\ldots,i_{2})\in[m+n]^{d-1}}L(e_{i_{d}})\cdots L(e_{i_{2}})\Bigg{(}\sum_{i_{1}=1}^{m+n}L(e_{i_{1}})L(e_{i_{1}}^{\ast})\Bigg{)}\big{(}L(e_{i_{2}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\big{)}$
$=\frac{1}{d!}\sum_{(i_{d},\ldots,i_{2})\in[m+n]^{d-1}}L(e_{i_{d}})\cdots L(e_{i_{2}})\big{(}L(e_{i_{2}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\big{)}$
$=\frac{1}{d!}\sum_{(i_{d},\ldots,i_{3})\in[m+n]^{d-2}}L(e_{i_{d}})\cdots L(e_{i_{3}})\Bigg{(}\sum_{i_{2}=1}^{m+n}L(e_{i_{2}})L(e_{i_{2}}^{\ast})\Bigg{)}\big{(}L(e_{i_{3}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\big{)}$
$=\frac{1}{d!}\sum_{(i_{d},\ldots,i_{3})\in[m+n]^{d-2}}L(e_{i_{d}})\cdots L(e_{i_{3}})\big{(}2L(e_{i_{3}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\big{)}$
=$\cdots$= ⋯
$=\frac{1}{d!}d!\varphi$
$=\varphi.$
4. Generalizations of the Schur-Sergeev duality
Let $W$ be a vector space and $A$ a subset of $\mathrm{End}_{\mathbb{C}}(W)$. The commutant of $A$ is denoted by $\mathrm{Comm}(A)$. That is
$\mathrm{Comm}(A)=\{f\in\mathrm{End}_{\mathbb{C}}(W)|fg=gf\text{ for all }g\in A\}.$
For any $d\geq 1$ and $0\leq k\leq d$, we consider $\mathrm{End}_{\mathbb{C}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$. Denote by $\mathcal{A}_{d}^{k}$ the subspace of $\mathrm{End}_{\mathbb{C}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$ spanned by operators of the form
$L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{\ast})\cdots L(v_{d}^{\ast}),$
with $w_{1},\ldots,w_{d}\in V$, $\sigma\in\mathfrak{S}_{d+k}$, and $v_{1}^{\ast},\ldots,v_{k}^{\ast}\in V^{\ast}$. For any $t\in\mathbb{C}\mathfrak{S}_{d+k}$, we define $\pi_{d+k}(t)\in\mathrm{End}_{\mathbb{C}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$ by $\pi_{d+k}(t)(\varphi)=\varphi\cdot t$ where $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$. Denote $\mathcal{B}_{d+k}=\pi_{d+k}(\mathbb{C}\mathfrak{S}_{d+k})$.
Theorem 4.1.
Keep the notations above. We have
$\mathrm{Comm}(\mathcal{A}_{d}^{k})=\mathcal{B}_{d+k},$
and
$\mathrm{Comm}(\mathcal{B}_{d+k})=\mathcal{A}_{d}^{k}.$
Proof.
Apparently, by Proposition 2.1, we have $\mathcal{A}_{d}^{k}\subset\mathrm{Comm}(\mathcal{B}_{d+k})$. Now for any $f\in\mathrm{Comm}(\mathcal{B}_{d+k})$ and $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$, it follows from Proposition 3.6 that
$f(\varphi)$
$=f(A_{d}(\varphi))$
$=f\Bigg{(}\frac{1}{d!}\sum_{(i_{d},\ldots,i_{1})\in[m+n]^{d}}L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\Bigg{)}$
$=f\Bigg{(}\frac{1}{d!}\sum_{(i_{d},\ldots,i_{1})\in[m+n]^{d}}e_{i_{d}}\cdots e_{i_{1}}\otimes L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\Bigg{)}$
$=f\Bigg{(}\frac{1}{d!}\sum_{(i_{d},\ldots,i_{1})\in[m+n]^{d}}\pi_{d+k}\big{(}L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\big{)}(e_{i_{d}}\cdots e_{i_{1}})\Bigg{)}$
$=\frac{1}{d!}\sum_{(i_{d},\ldots,i_{1})\in[m+n]^{d}}\pi_{d+k}\big{(}L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\big{)}f(e_{i_{d}}\cdots e_{i_{1}})$
$=\frac{1}{d!}\sum_{(i_{d},\ldots,i_{1})\in[m+n]^{d}}f(e_{i_{d}}\cdots e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi).$
Write $f(e_{i_{d}}\cdots e_{i_{1}})=\sum_{J=(j_{d},\ldots,j_{1})\in[m+n]^{d}}e_{j_{d}}\cdots e_{j_{1}}\otimes t_{J}$ with $t_{J}\in\mathbb{C}\mathfrak{S}_{d+k}$. Then
$f(e_{i_{d}}\cdots e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)$
$=\sum_{J=(j_{d},\ldots,j_{1})\in[m+n]^{d}}L(e_{j_{d}})\cdots L(e_{j_{1}})L(t_{J})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi).$
Hence $f\in\mathcal{A}_{d}^{k}$. As a consequence, $\mathrm{Comm}(\mathcal{B}_{d+k})\subset\mathcal{A}_{d}^{k}$.
Since $\mathbb{C}\mathfrak{S}_{d+k}$ is a semisimple algebra, it follows from the double commutant theorem that $\mathrm{Comm}(\mathcal{A}_{d}^{k})=\mathcal{B}_{d+k}$.∎
As an application of the above theorem, we provide a new proof of the Schur-Sergeev duality theorem. Firstly, we need the following technical lemma.
Lemma 4.2.
For any $i,j,k\in[m+n]$, we have
$L(e_{i})L(e_{j})L(e_{k}^{\ast})=(-1)^{|e_{i}|(|e_{j}|+|e_{k}^{\ast}|)}\big{(}L(e_{j})L(e_{k}^{\ast})L(e_{i})-\delta_{ik}L(e_{j})\big{)},$
on $\overline{T}_{d}$.
Proof.
For any $(l_{d},\ldots,l_{1})\in[m+n]^{d}$ and $\sigma\in\mathfrak{S}_{\infty}$, we have
$(-1)^{|e_{i}|(|e_{j}|+|e_{k}^{\ast}|)}\big{(}L(e_{j})L(e_{k}^{\ast})L(e_{i})-\delta_{ik}L(e_{j})\big{)}(e_{l_{d}}\cdots e_{l_{1}}\otimes\sigma)$
$=(-1)^{|e_{i}|(|e_{j}|+|e_{k}^{\ast}|)}\big{(}\rho_{d}(E_{jk})(e_{i}e_{l_{d}}\cdots e_{l_{1}})\otimes\sigma-\delta_{ik}e_{j}e_{l_{d}}\cdots e_{l_{1}}\otimes\sigma\big{)}$
$=(-1)^{|e_{i}|(|e_{j}|+|e_{k}^{\ast}|)}\big{(}(E_{jk}e_{i})e_{l_{d}}\cdots e_{l_{1}}\otimes\sigma-\delta_{ik}e_{j}e_{l_{d}}\cdots e_{l_{1}}\otimes\sigma\big{)}$
$\ \ \ \ +(-1)^{|e_{i}|(|e_{j}|+|e_{k}^{\ast}|)}\sum_{p=1}^{d}(-1)^{|E_{jk}|(|e_{i}|+|e_{l_{d}}|+\cdots+|e_{l_{p+1}}|)}e_{i}e_{l_{d}}\cdots(E_{jk}e_{l_{p}})\cdots e_{l_{1}}\otimes\sigma$
$=\sum_{p=1}^{d}(-1)^{|E_{jk}|(|e_{l_{d}}|+\cdots+|e_{l_{p+1}}|)}e_{i}e_{l_{d}}\cdots(E_{jk}e_{l_{p}})\cdots e_{l_{1}}\otimes\sigma$
$=L(e_{i})\Bigg{(}\sum_{p=1}^{d}(-1)^{|E_{jk}|(|e_{l_{d}}|+\cdots+|e_{l_{p+1}}|)}e_{l_{d}}\cdots(E_{jk}e_{l_{p}})\cdots e_{l_{1}}\otimes\sigma\Bigg{)}$
$=L(e_{i})(\rho_{d}(E_{jk})(e_{l_{d}}\cdots e_{l_{1}})\otimes\sigma)$
$=L(e_{i})L(e_{j})L(e_{k}^{\ast})(e_{l_{d}}\cdots e_{l_{1}}\otimes\sigma),$
as desired.∎
Theorem 4.3 (Schur-Sergeev duality).
For each $d\geq 1$, we have
$\mathrm{Comm}\big{(}\rho_{d}(U(gl(m|n)))\big{)}=\mathcal{B}_{d},$
and
$\mathrm{Comm}(\mathcal{B}_{d})=\rho_{d}(U(gl(m|n))).$
Proof.
We are going to show that $\mathcal{A}_{d}^{0}=\rho_{d}(U(gl(m|n)))$. We first notice that every element in $\mathcal{A}_{d}^{0}$ is a linear combination of operators of the form
$L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{\ast})\cdots L(v_{d}^{\ast}),$
with $w_{1},\ldots,w_{d}\in V$, $\sigma\in\mathfrak{S}_{d}$, and $v_{1}^{\ast},\ldots,v_{k}^{\ast}\in V^{\ast}$. By Proposition 3.1, the above operators can be rewritten in the form
$\eta L(w^{\prime}_{d})\cdots L(w^{\prime}_{1})L(v_{1}^{\ast})\cdots L(v_{d}^{\ast}),$
for some number $\eta$ and $w^{\prime}_{1},\ldots,w^{\prime}_{d}\in V$. We identify $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d}$ with $V^{\otimes d}$ in a natural way.
For any $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d}$, by (3.1), we have
$\rho_{d}(E_{ij})(\varphi)$ $=L(e_{i})L(e_{j}^{\ast})(\varphi)$
$=L(e_{i})A_{d-1}L(e_{j}^{\ast})(\varphi)$
$=\frac{1}{(d-1)!}\sum_{k_{1},\ldots,k_{d-1}=1}^{m+n}L(e_{i})L(e_{k_{d-1}})\cdots L(e_{k_{1}})L(e_{k_{1}}^{\ast})\cdots L(e_{k_{d-1}}^{\ast})L(e_{j}^{\ast})(\varphi).$
Therefore $\rho_{d}(E_{ij})\in\mathcal{A}_{d}^{0}$, and hence $\rho_{d}(U(gl(m|n)))\subset\mathcal{A}_{d}^{0}$.
On the other hand, for any $(i_{p},\ldots,i_{1}),(j_{p},\ldots,j_{1})\in[m+n]^{p}$, we use induction on $p$ to show that $L(e_{i_{p}})\cdots L(e_{i_{1}})L(e_{j_{1}}^{\ast})\cdots L(e_{j_{p}}^{\ast})\in\rho_{d}(U(gl(m|n)))$. The case $p=1$ follows from (3.1). Assume that the result holds for $p\geq 1$. By the above lemma, it is easy to see that
$L(e_{i_{p+1}})\cdots L(e_{i_{1}})L(e_{j_{1}}^{\ast})\cdots L(e_{j_{p+1}}^{\ast})$
$=\kappa L(e_{i_{1}})L(e_{j_{1}}^{\ast})L(e_{i_{p+1}})\cdots L(e_{i_{2}})L(e_{j_{2}}^{\ast})\cdots L(e_{j_{p+1}}^{\ast})+\text{terms in }\mathcal{A}_{d-2}^{0}$
$=\kappa\rho_{d}(E_{i_{1}j_{1}})L(e_{i_{p+1}})\cdots L(e_{i_{2}})L(e_{j_{2}}^{\ast})\cdots L(e_{j_{p+1}}^{\ast})+\text{terms in }\mathcal{A}_{d-2}^{0},$
for some number $\kappa$. By the inductive hypothesis, we get the result. ∎
5. Extensions through Hecke algebras
5.1. Hecke algebras
Let $q$ be an indeterminate. The Hecke algebra $\mathscr{H}_{d}$ of type $A_{d-1}$ is a $\mathbb{C}(q)$-algebra generated by $T_{1},T_{2},\ldots,T_{d-1}$ subject to
$(T_{i}-q)(T_{i}+q^{-1})=0,$
$T_{i}T_{i+1}T_{i}=T_{i+1}T_{i}T_{i+1},$
$T_{i}T_{j}=T_{j}T_{i},\qquad\mbox{for $|i-j|>1$}.$
For any reduced expression $s_{i_{1}}s_{i_{2}}\cdots s_{i_{k}}$ of $w\in\mathfrak{S}_{d}$, we denote $T_{w}=T_{i_{1}}T_{i_{2}}\cdots T_{i_{k}}$. It is well-known that $T_{w}$ is independent of the choice of reduced expressions of $w$.
Let $\mathscr{H}_{\infty}$ be the inductive limit of the natural inclusions $\mathscr{H}_{0}\subset\mathscr{H}_{1}\subset\cdots$. Given $k\geq 1$, we define an algebra homomorphism :${}^{\uparrow k}:\mathscr{H}_{\infty}\rightarrow\mathscr{H}_{\infty}$ by $T_{i}^{\uparrow k}=T_{i+k}$.
Define a map $\gamma:[m+n]\times[m+n]\to\{-1,1\}$ by
$\gamma(i,j)=\left\{\begin{array}[]{ll}1,&\mbox{if $i>j$},\\ -1,&\mbox{if $i\leq j$}.\end{array}\right.$
The following lemma will be used in the next section.
Lemma 5.1.
For any $1\leq i,j,k\leq m+n$, we have
$T_{1}^{\gamma(i,j)}T_{2}^{\gamma(i,k)}T_{1}^{\gamma(j,k)}=T_{2}^{\gamma(j,k)}T_{1}^{\gamma(i,k)}T_{2}^{\gamma(i,j)}.$
Proof.
It can be verified by the braid relation $T_{1}T_{2}T_{1}=T_{2}T_{1}T_{2}$ case by case. ∎
5.2. $q$-tensor superspaces
We introduce a parity on $[m+n]$ by
$\widehat{i}=\left\{\begin{array}[]{ll}0,&\mbox{if $1\leq i\leq m$},\\ 1,&\mbox{if $m<i\leq m+n$}\end{array}\right.$
and let $q_{i}=q^{(-1)^{\widehat{i}}}$.
Let $\mathbb{V}=\mathbb{V}_{\overline{0}}\oplus\mathbb{V}_{\overline{1}}$ be a vector space over $\mathbb{C}(q)$ with a specified basis $\{e_{1},e_{2},\ldots,e_{m+n}\}$, where $\mathbb{V}_{\overline{0}}$ (resp. $\mathbb{V}_{\overline{1}}$) is spanned by $\{e_{i}|1\leq i\leq m\}$ (resp. $\{e_{i}|m+1\leq i\leq m+n\}$).
For $I=(i_{d},\ldots,i_{2},i_{1})\in[m+n]^{d}$, we denote
$e_{I}=e_{i_{d}}\cdots e_{i_{2}}e_{i_{1}}.$
Then $\{e_{I}~{}|~{}I\in[m+n]^{d}\}$ forms a basis of $\mathbb{V}^{\otimes d}$.
As before, $\mathfrak{S}_{d}$ acts on $[m+n]^{d}$ from the right hand side. The right $\mathscr{H}_{d}$-action on $\mathbb{V}^{\otimes d}$ is defined on the basis element $e_{I}$ as follows: for $1\leq k<d$,
(5.1) $e_{I}.T_{k}=\left\{\begin{array}[]{ll}(-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}},&\mbox{if $i_{k}>i_{k+1}$},\\ (-1)^{\widehat{i_{k}}}q_{i_{k}}e_{I},&\mbox{if $i_{k}=i_{k+1}$},\\ (-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}}+(q-q^{-1})e_{I},&\mbox{if $i_{k}<i_{k+1}$}.\end{array}\right.$
We can rewrite (5.1) as
(5.2) $e_{I}.T_{k}^{\gamma(i_{k},i_{k+1})}=(-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}q_{i_{k}}^{-\delta_{i_{k}i_{k+1}}}e_{Is_{k}},$
for $1\leq k<d$. Since $\gamma(j,i)=-2\delta_{i,j}-\gamma(i,j)$, we have
$e_{I}.T_{k}^{\gamma(i_{k+1},i_{k})}=e_{I}.(T_{k}^{-2\delta_{i_{k}i_{k+1}}}T_{k}^{-\gamma(i_{k},i_{k+1})})=q_{i_{k}}^{-2\delta_{i_{k}i_{k+1}}}e_{I}.T_{k}^{-\gamma(i_{k},i_{k+1})},$
for $1\leq k<d$. The above equation together with the fact $T_{k}^{-1}=T_{k}+(q^{-1}-q)$ implies
$e_{I}.T_{k}^{\gamma(i_{k+1},i_{k})}=e_{I}.(T_{k}^{-\gamma(i_{k},i_{k+1})}-q^{\delta_{i_{k},i_{k+1}}}+q^{-\delta_{i_{k},i_{k+1}}}),$
for $1\leq k<d$.
5.3. Super-extensions
We set
$\widetilde{T}_{d}(\mathbb{V})=\mathrm{Ind}_{\mathscr{H}_{d}}^{\mathscr{H}_{\infty}}\mathbb{V}^{\otimes d}=\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{\infty},$
and
$\widetilde{T}(\mathbb{V})=\bigoplus_{d\geq 0}\widetilde{T}_{d}(\mathbb{V}).$
Endow an algebra structure on $\widehat{T}(\mathbb{V})$ whose product $\cdot$ is defined by
$(e_{j_{k}}\cdots e_{j_{1}}\otimes T_{\tau})\cdot(e_{i_{d}}\cdots e_{i_{1}}\otimes T_{\sigma})=e_{j_{k}}\cdots e_{j_{1}}e_{i_{d}}\cdots e_{i_{1}}\otimes T_{\tau}^{\uparrow d}T_{\sigma}.$
The product $\varphi\cdot T$, with $\varphi\in\widetilde{T}(\mathbb{V})$ and $T\in\mathscr{H}_{\infty}$, provides a natural right $\mathscr{H}_{\infty}$-action on $\widetilde{T}(\mathbb{V})$.
5.4. Multiplications and Derivations
For any $\varphi\in\widetilde{T}(\mathbb{V})$, define
$L(\varphi):\widetilde{T}(\mathbb{V})$$\longrightarrow$ $\widetilde{T}(\mathbb{V}),$
[…]
For $j\in[m+n]$, define $L(e_{j}^{*}):\widetilde{T}(\mathbb{V})\rightarrow\widetilde{T}(\mathbb{V})$ by
(5.3) $L(e_{j}^{*})(e_{i_{d}}\cdots e_{i_{1}}\otimes T_{\sigma})$
$=$ $\sum_{k=1}^{d}(-1)^{\widehat{j}(\widehat{i_{d}}+\cdots+\widehat{i_{k+1}})}\langle e_{j}^{*},e_{i_{k}}\rangle g_{j}(e_{i_{d}})\cdots g_{j}(e_{i_{k+1}})f_{j}(e_{i_{k-1}})\cdots f_{j}(e_{i_{1}})\otimes T_{\sigma},$
where $f_{j}:\mathbb{V}\rightarrow\mathbb{V}$ is the linear endomorphism defined by
$f_{j}(e_{r})={q_{j}}^{-\delta_{jr}}e_{r}$
and $g_{j}:\mathbb{V}\rightarrow\mathbb{V}\otimes\mathscr{H}_{2}$ is the linear map defined by
$g_{j}(e_{r})=e_{r}T_{1}^{-\gamma(j,r)}.$
It is easy to see that $L(\varphi)$ is well-defined. For $L(e_{j}^{*})$, we only need to show that it commutes with the action of $\mathscr{H}_{\infty}$. We introduce a linear map $h_{k}^{(j)}:T_{d}(\mathbb{V})\rightarrow\widetilde{T}_{d}(\mathbb{V})$ for $1\leq k\leq d$ defined by
$h_{k}^{(j)}(e_{i_{d}}\cdots e_{i_{1}})=(-1)^{\widehat{j}(\widehat{i_{d}}+\cdots+\widehat{i_{k+1}})}\langle e_{j}^{*},e_{i_{k}}\rangle g_{j}(e_{i_{d}})\cdots g_{j}(e_{i_{k+1}})f_{j}(e_{i_{k-1}})\cdots f_{j}(e_{i_{1}}).$
It is just the $k$-th term of $L(e_{j}^{*})$ over $T_{d}(\mathbb{V})$ in (5.3). For any $T\in\mathscr{H}_{\infty}$, we define $\pi(T)\in\mathrm{End}_{\mathbb{C}}(\widetilde{T}(\mathbb{V}))$ by $\pi(T)(\varphi)=\varphi\cdot T$ for $\varphi\in\widetilde{T}(\mathbb{V})$. To prove the well-definedness of $L(e_{j}^{*})$, it suffices to show that $\pi(T_{1}),...,\pi(T_{d-1})$ commutes with $\sum_{k=1}^{d}h_{k}^{(j)}$. We have the following lemma:
Lemma 5.2.
Write $h_{k}$ for $h_{k}^{(j)}$. Then for $r=1,...,d-1$, we have
  1. (i)
    $h_{k}$ commutes with $\pi(T_{r})$ unless $k=r,r+1$,
  2. (ii)
    $h_{r+1}+h_{r}$ commutes with $\pi(T_{r})$.
Proof.
Fix $i_{1},...,i_{d}$. By the definition of $h_{k}$, we have the following relation between $h_{k}$ and $h_{k+1}$:
(5.4) $h_{k}=q_{i}^{2\delta_{j,i_{k+1}}}\pi(T_{k}^{-\gamma(i,i_{k+1})})h_{k+1}\pi(T_{k}^{\gamma(i_{k},i_{k+1})}).$
Then when $r=k+1$ and $d\geq 3$, we have
$h_{k}\pi(T_{k+1}^{\gamma(i_{k+1},i_{k+2})})$
$=q_{i}^{2(\delta_{j,i_{k+1}}+\delta_{j,i_{k+2}})}T_{k}^{-\gamma(i_{k},i_{k+2})}\pi(T_{k+1}^{-\gamma(i_{k},i_{k+1})})h_{k+2}\pi(T_{k+1}^{\gamma(i_{k},i_{k+1})}T_{k}^{\gamma(i_{k},i_{k+2})}T_{k+1}^{\gamma(i_{k+1},i_{k+2})})\delta_{j,i_{k}}$
$=q_{i}^{2(\delta_{j,i_{k+1}}+\delta_{j,i_{k+2}})}\pi(T_{k+1}^{\gamma(i_{k+1},i_{k+2})}T_{k}^{-\gamma(i_{k},i_{k+1})}T_{k+1}^{-\gamma(i_{k},i_{k+2})})h_{k+2}\pi(T_{k+1}^{\gamma(i_{k},i_{k+2})}T_{k}^{\gamma(i_{k},i_{k+1})})\delta_{j,i_{k}}$
$=\pi(T_{k+1}^{\gamma(i_{k+1},i_{k+2})})h_{k},$
where the first equality follows from (5.4) and the second one from Lemma 5.1.
When $k+2\leq r\leq d-1$, we show that $\pi(T_{r})$ commutes with $h_{k}$ by induction on $k$. We start the induction by $\pi(T_{r})h_{r-1}=h_{r-1}\pi(T_{r})$ which is the result of the last paragraph. For $k<r-1$, we have
$\pi(T_{r})h_{k}$ $=\pi(T_{r})q_{j}^{2\gamma(j,i_{k+1})}\pi(T_{k}^{-\gamma(j,i_{k+1})})h_{k+1}\pi(T_{k}^{\gamma(i_{k},i_{k+1})})\delta_{j,i_{k}}$
$=q_{j}^{2\delta_{j,i_{k+1}}}\pi(T_{k}^{-\gamma(j,i_{k+1})}T_{r})h_{k+1}\pi(T_{k}^{\gamma(i_{k},i_{k+1})})\delta_{j,i_{k}}$
$=q_{j}^{2\delta_{j,i_{k+1}}}\pi(T_{k}^{-\gamma(j,i_{k+1})})h_{k+1}\pi(T_{k}^{\gamma(i_{k},i_{k+1})})\pi(T_{r})\delta_{j,i_{k}}$
$=h_{k}\pi(T_{r}),$
which proves (i).
For (ii), notice that
$h_{k+1}\pi(T_{k}^{\gamma(i_{k},i_{k+1})})=q_{j}^{-2\delta_{j,i_{k+1}}}\pi(T_{k}^{\gamma(j,i_{k+1})})h_{k},$
and
$h_{k}\pi(T_{k}^{\gamma(i_{k},i_{k+1})})$ $=q_{j}^{2\delta_{j,i_{k}}}\pi(T_{k}^{-\gamma(i_{k+1},i_{k})})h_{k+1}\pi(T_{k}^{\gamma(i_{k+1},i_{k})}T_{k}^{\gamma(i_{k},i_{k+1})})$
$=q_{j}^{2\delta_{j,i_{k}}}\pi(T_{k}^{-\gamma(i_{k+1},i_{k})})h_{k+1}\pi(T_{k}^{-2\delta_{i_{k},i_{k+1}}})$
$=q_{j}^{2\delta_{j,i_{k}}}\pi(T_{k}^{-\gamma(i_{k+1},i_{k})})h_{k+1}{q_{j}}^{-2\delta_{i_{k},i_{k+1}}}$
$=\pi(T_{k}^{-\gamma(i_{k+1},i_{k})})h_{k+1}$
$=\pi(T_{k}^{\gamma(i_{k},i_{k+1})}+q^{\delta_{i_{k},i_{k+1}}}-q^{-\delta_{i_{k},i_{k+1}}})h_{k+1}$
$=\pi(T_{k}^{\gamma(i_{k},i_{k+1})})h_{k+1}+(q^{\delta_{i_{k},i_{k+1}}}-q^{-\delta_{i_{k},i_{k+1}}})h_{k+1}$
$=\pi(T_{k}^{\gamma(i_{k},i_{k+1})})h_{k+1}+(-1)^{\widehat{i}}(q^{\delta_{i_{k},i_{k+1}}}-q^{-\delta_{i_{k},i_{k+1}}})\pi(T_{k}^{\gamma(j,i_{k+1})})q_{j}^{-\delta_{j,i_{k+1}}}h_{k}$
$=\pi(T_{k}^{\gamma(i_{k},i_{k+1})})h_{k+1}+(1-q_{j}^{-2\delta_{j,i_{k+1}}})\pi(T_{k}^{\gamma(j,i_{k+1})})h_{k}$
$=\pi(T_{k}^{\gamma(i_{k},i_{k+1})})h_{k+1}+\pi(T_{k}^{\gamma(j,i_{k+1})})h_{k}-q_{j}^{-2\delta_{j,i_{k+1}}}\pi(T_{k}^{\gamma(j,i_{k+1})})h_{k}.$
Adding the above two equations, we get the result. ∎
5.5. General linear quantum supergroups
The general linear quantum supergroup $\mathbf{U}_{m|n}=U_{q}(\mathfrak{gl}_{m|n})$ is a $\mathbb{C}(q)$-algebra generated by
$\left\{\begin{array}[]{lll}\mbox{even generators:}&E_{i},F_{i},K_{j}^{\pm 1},&1\leq i,j\leq m+n,i\neq m,m+n;\\ \mbox{odd generators:}&E_{m},F_{m}&\end{array}\right.$
subject to the following relations:
$K_{i}K_{j}=K_{j}K_{i},\quad K_{i}K_{i}^{-1}=K_{i}^{-1}K_{i}=1;$
$K_{i}E_{j}=q_{i}^{\delta_{ij}-\delta_{i,j+1}}E_{j}K_{i},\quad K_{i}F_{j}=q_{i}^{\delta_{i,j+1}-\delta_{ij}}F_{j}K_{i};$
$E_{i}F_{j}-(-1)^{\delta_{im}\delta_{jm}}F_{j}E_{i}=\delta_{ij}\frac{K_{i}K_{i+1}^{-1}-K_{i}^{-1}K_{i+1}}{q_{i}-q_{i}^{-1}};$
$E_{i}E_{j}=E_{j}E_{i},\quad F_{i}F_{j}=F_{j}F_{i},\quad(|i-j|>1);$
$E_{i}^{2}E_{j}-(q_{i}+q_{i}^{-1})E_{i}E_{j}E_{i}+E_{j}E_{i}^{2}=0,\quad(i\neq m,|i-j|=1);$
$F_{i}^{2}F_{j}-(q_{i}+q_{i}^{-1})F_{i}F_{j}F_{i}+F_{j}F_{i}^{2}=0,\quad(i\neq m,|i-j|=1);$
$E_{m}^{2}=F_{m}^{2}=E_{m}E_{m-1,m+2}+E_{m-1,m+2}E_{m}=F_{m}F_{m-1,m+2}+F_{m-1,m+2}F_{m}=0,$
where
$E_{m-1,m+2}$ $=E_{m-1}E_{m}E_{m+1}-q^{-1}E_{m-1}E_{m+1}E_{m}-qE_{m}E_{m+1}E_{m-1}+E_{m+1}E_{m}E_{m-1},$
$F_{m-1,m+2}$ $=F_{m-1}F_{m}F_{m+1}-q^{-1}F_{m-1}F_{m+1}F_{m}-qF_{m}F_{m+1}F_{m-1}+F_{m+1}F_{m}F_{m-1}.$
There is a Hopf superalgebra structure on $\mathbf{U}_{m|n}$, whose comultiplication $\Delta$ is
$\Delta(K_{i})=K_{i}\otimes K_{i},$
$\Delta(E_{i})=1\otimes E_{i}+E_{i}\otimes K_{i}K_{i+1}^{-1},$
$\Delta(F_{i})=F_{i}\otimes 1+K_{i+1}K_{i}^{-1}\otimes F_{i}.$
Since $\Delta$ is coassociative, we can define
$\Delta^{(d)}=(\Delta\otimes\mathrm{id}^{\otimes d-1})\Delta^{(d-1)},$
for $d\geq 2$. Here we set $\Delta^{(1)}=\Delta$ for convenience.
The following lemma can be derived by a straightforward computation.
Lemma 5.3.
For $1\leq i\leq m+n$, we have
$\Delta^{(d)}(K_{i})$ $=K_{i}^{\otimes d+1},$
$\Delta^{(d)}(E_{i})$ $=1\otimes\Delta^{(d-1)}(E_{i})+E_{i}\otimes(K_{i}K_{i+1}^{-1})^{\otimes d},$
$\Delta^{(d)}(F_{i})$ $=F_{i}\otimes 1^{\otimes d}+K_{i+1}K_{i}^{-1}\otimes\Delta^{(d-1)}(F_{i}).$
Particularly, $|\Delta^{(d)}(E_{i})|=|E_{i}|$ and $|\Delta^{(d)}(F_{i})|=|F_{i}|$, where ||$\cdot$|| ⋅ | means the parity of elements in $\mathbf{U}_{m|n}$.
The superspace $\mathbb{V}$ admits a natural representation of $\mathbf{U}_{m|n}$ via
$K_{i}.e_{j}=q_{i}^{\delta_{ij}}e_{j},\quad E_{i}.e_{j}=\delta_{i+1,j}e_{i},\quad F_{i}.e_{j}=\delta_{i,j}e_{i+1}.$
Then the $q$-tensor superspace $\mathbb{V}^{\otimes d}$ is also equipped with a $\mathbf{U}_{m|n}$-module structure via $\Delta^{(d)}$. We shall show that the $\mathbf{U}_{m|n}$-module structure can be extended over the $\widetilde{T}(\mathbb{V})$ as $x.(e_{I}\otimes T_{\sigma})=(x.e_{I})\otimes T_{\sigma}$ for $x\in\mathbf{U}_{m|n},e_{I}\in\mathbb{V}^{\otimes d},T_{\sigma}\in\mathscr{H}_{\infty}$. It can be seen instantly for $K$, since
$K_{j}.(e_{i_{d}}\cdots e_{i_{1}})=(q_{j})^{\delta_{ji_{d}}+\cdots+\delta_{ji_{1}}}e_{i_{d}}\cdots e_{i_{1}},$
which commutes with $\mathscr{H}_{\infty}$. The commutativity of $E$ and $F$ with $\mathscr{H}_{\infty}$ will be proved later (see 6.7).
6. Properties of operators in quantum version
6.1. Commutation relations
Proposition 6.1.
For any $i,j\in[m+n]$ and $1\leq k<d$, as operators on $\widetilde{T}(\mathbb{V})$, we have
$K_{j}L(e_{i})=q_{j}^{\delta_{ij}}L(e_{i})K_{j},\quad K_{j}L(e_{i}^{*})=q_{j}^{-\delta_{ij}}L(e_{i}^{*})K_{j},\quad K_{j}L(T_{k})=L(T_{k})K_{j}$
Proof.
Check directly. ∎
Lemma 6.2.
For $1\leq j\leq m+n$,
$L(e_{j})L(e_{j}^{*})=\frac{K_{j}-K_{j}^{-1}}{q_{j}-q_{j}^{-1}}.$
Proof.
For $I=(i_{d},\ldots,i_{2},i_{1})\in[m+n]^{d}$, by a direct computation, we have
$\phantom{=}L(e_{j})L(e_{j}^{*})(e_{i_{d}}\cdots e_{i_{1}}\otimes T_{\sigma})$
$=\sum_{k=1}^{d}(-1)^{\widehat{j}(\widehat{i_{d}}+\cdots+\widehat{i_{k+1}})}\langle e_{j}^{*},e_{i_{k}}\rangle e_{j}g_{j}(e_{i_{d}})\cdots g_{j}(e_{i_{k+1}})f_{j}(e_{i_{k-1}})\cdots f_{j}(e_{i_{1}})\otimes T_{\sigma}$
$=\sum_{k=1}^{d}\langle e_{j}^{*},e_{i_{k}}\rangle e_{j}e_{i_{d}}(-1)^{\widehat{j}\widehat{i_{d}}}T_{1}^{-\gamma(j,i_{d})}\cdots e_{i_{k+1}}(-1)^{\widehat{j}\widehat{i_{k+1}}}T_{1}^{-\gamma(j,i_{k+1})}f_{j}(e_{i_{k-1}})\cdots f_{j}(e_{i_{1}})\otimes T_{\sigma}$
[…]
$=\sum_{k=1}^{d}q_{j}^{(\delta_{ji_{d}}+\cdots+\delta_{ji_{k+1}})-(\delta_{ji_{k-1}}+\cdots+\delta_{ji_{1}})}\langle e_{j}^{*},e_{i_{k}}\rangle e_{i_{d}}\cdots e_{i_{k+1}}e_{i_{k}}e_{i_{k-1}}\cdots e_{i_{1}}\otimes T_{\sigma}$
$=\Bigg{(}\sum_{k=1}^{d}\delta_{ji_{k}}q_{j}^{(\delta_{ji_{d}}+\cdots+\delta_{ji_{k+1}})-(\delta_{ji_{k-1}}+\cdots+\delta_{ji_{1}})}\Bigg{)}e_{i_{d}}\cdots e_{i_{k+1}}e_{i_{k}}e_{i_{k-1}}\cdots e_{i_{1}}\otimes T_{\sigma}.$
Note that
$\sum_{k=1}^{d}\delta_{ji_{k}}q_{j}^{(\delta_{ji_{d}}+\cdots+\delta_{ji_{k+1}})-(\delta_{ji_{k-1}}+\cdots+\delta_{ji_{1}})}$
$=\sum_{k=1}^{d}\frac{q_{j}^{(\delta_{ji_{d}}+\cdots+\delta_{ji_{k+1}}+\delta_{ji_{k}})-(\delta_{ji_{k-1}}+\cdots+\delta_{ji_{1}})}-q_{j}^{(\delta_{ji_{d}}+\cdots+\delta_{ji_{k+1}})-(\delta_{ji_{k}}+\delta_{ji_{k-1}}+\cdots+\delta_{ji_{1}})}}{q_{j}-q_{j}^{-1}}$
$=\frac{q_{j}^{(\delta_{ji_{d}}+\cdots+\delta_{ji_{1}})}-q_{j}^{-(\delta_{ji_{d}}+\cdots+\delta_{ji_{1}})}}{q_{j}-q_{j}^{-1}}.$
We get the desired result. ∎
Proposition 6.3.
For any $1\leq i,j\leq m+n$, we have
(6.1) $L(e_{i})L(e_{j})$ $=(-1)^{\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(e_{j})L(e_{i})L(T_{1}^{\gamma(i,j)}),$
(6.2) $L(e_{i}^{*})L(e_{j}^{*})$ $=(-1)^{\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(T_{1}^{\gamma(i,j)})L(e_{j}^{*})L(e_{i}^{*}),$
(6.3) $L(e_{i}^{*})L(e_{j})$ $=\delta_{ij}K_{i}^{-1}+(-1)^{\widehat{i}\widehat{j}}L(e_{j})L(T_{1}^{-\gamma(i,j)})L(e_{i}^{*}).$
Proof.
The relation (6.1) follows from
$L(e_{i})L(e_{j})(e_{i_{d}}\cdots e_{i_{1}})$ $=e_{i}e_{j}e_{i_{d}}\cdots e_{i_{1}}$
$=(-1)^{\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}e_{j}e_{i}T_{1}^{\gamma(i,j)}e_{i_{d}}\cdots e_{i_{1}}$
$=(-1)^{\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(e_{j})L(e_{i})L(T_{1}^{\gamma(i,j)})e_{i_{d}}\cdots e_{i_{1}},$
while (6.3) from
$L(e_{i}^{*})L(e_{j})(e_{i_{d}}\cdots e_{i_{1}})$
$=L(e_{i}^{*})(e_{j}e_{i_{d}}\cdots e_{i_{1}})$
$=\langle e_{i}^{*},e_{j}\rangle q_{i}^{-\delta_{ii_{d}}-\cdots-\delta_{ii_{1}}}e_{i_{d}}\cdots e_{i_{1}}$
$\qquad+\sum_{k=1}^{d}(-1)^{\widehat{i}(\widehat{j}+\widehat{i_{d}}+\cdots+\widehat{i_{k+1}})}\langle e_{j}^{*},e_{i_{k}}\rangle g_{i}(e_{j})g_{i}(e_{i_{d}})\cdots g_{i}(e_{i_{k+1}})f_{i}(e_{i_{k-1}})\cdots f_{i}(e_{i_{1}})$
$=\langle e_{i}^{*},e_{j}\rangle K_{i}^{-1}.e_{i_{d}}\cdots e_{i_{1}}$
$\qquad+\sum_{k=1}^{d}(-1)^{\widehat{i}(\widehat{j}+\widehat{i_{d}}+\cdots+\widehat{i_{k+1}})}\langle e_{j}^{*},e_{i_{k}}\rangle g_{i}(e_{j})g_{i}(e_{i_{d}})\cdots g_{i}(e_{i_{k+1}})f_{i}(e_{i_{k-1}})\cdots f_{i}(e_{i_{1}})$
$=\big{(}\delta_{ij}K_{i}^{-1}+(-1)^{\widehat{i}\widehat{j}}L(e_{j})L(T_{1}^{-\gamma(i,j)})L(e_{i}^{*})\big{)}(e_{i_{d}}\cdots e_{i_{1}}).$
Now let us prove (6.2), which is equivalent to
[…]
for $d=0,1,2,\ldots$. We shall prove (6.4) by induction on $d$. It is trivial for $d=0,1$. Take any $e_{i_{d}}\cdots e_{i_{1}}\otimes T_{\sigma}\in\widetilde{T}_{d}(\mathbb{V})$. If $i_{d}=k\neq i,j$, then we have
$L(e_{i}^{*})L(e_{j}^{*})(e_{i_{d}}\cdots e_{i_{1}}\otimes T_{\sigma})$
$=L(e_{i}^{*})L(e_{j}^{*})L(e_{k})(e_{i_{d-1}}\cdots e_{i_{1}}\otimes T_{\sigma})$
$=L(e_{i}^{*})(-1)^{\widehat{j}\widehat{k}}L(e_{k})L(T_{1}^{-\gamma(j,k)})L(e_{j}^{*})(e_{i_{d-1}}\cdots e_{i_{1}}\otimes T_{\sigma})$
$=(-1)^{(\widehat{i}+\widehat{j})\widehat{k}}L(e_{k})L(T_{1}^{-\gamma(i,k)})L(e_{i}^{*})L(T_{1}^{-\gamma(j,k)})L(e_{j}^{*})(e_{i_{d-1}}\cdots e_{i_{1}}\otimes T_{\sigma})$
$=(-1)^{(\widehat{i}+\widehat{j})\widehat{k}}L(e_{k})L(T_{1}^{-\gamma(i,k)}T_{2}^{-\gamma(j,k)})L(e_{i}^{*})L(e_{j}^{*})(e_{i_{d-1}}\cdots e_{i_{1}}\otimes T_{\sigma})$
$=(-1)^{(\widehat{i}+\widehat{j})\widehat{k}+\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(e_{k})L(T_{1}^{-\gamma(i,k)}T_{2}^{-\gamma(j,k)}T_{1}^{\gamma(i,j)})L(e_{j}^{*})L(e_{i}^{*})(e_{i_{d-1}}\cdots e_{i_{1}}\otimes T_{\sigma})$
[…]
$=(-1)^{(\widehat{i}+\widehat{j})\widehat{k}+\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(T_{1}^{\gamma(i,j)})L(e_{k})L(T_{1}^{-\gamma(j,k)}T_{2}^{-\gamma(i,k)})L(e_{j}^{*})L(e_{i}^{*})(e_{i_{d-1}}\cdots e_{i_{1}}\otimes T_{\sigma})$
$=(-1)^{\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(T_{1}^{\gamma(i,j)})L(e_{j}^{*})L(e_{i}^{*})L(e_{k})(e_{i_{d-1}}\cdots e_{i_{1}}\otimes T_{\sigma}).$
If $i_{d}=i$, then
$L(e_{i}^{*})L(e_{j}^{*})(e_{i_{d}}\cdots e_{i_{1}})$
$=L(e_{i}^{*})L(e_{j}^{*})L(e_{i})(e_{i_{d-1}}\cdots e_{i_{1}})$
$=L(e_{i}^{*})(\delta_{ij}K_{j}^{-1}+(-1)^{\widehat{i}\widehat{j}}L(e_{i})L(T_{1}^{-\gamma(j,i)})L(e_{j}^{*}))(e_{i_{d-1}}\cdots e_{i_{1}})$
$=(\delta_{ij}L(e_{i}^{*})K_{j}^{-1}+(-1)^{\widehat{i}\widehat{j}}L(e_{i}^{*})L(e_{i})L(T_{1}^{-\gamma(j,i)})L(e_{j}^{*}))(e_{i_{d-1}}\cdots e_{i_{1}})$
$=(\delta_{ij}L(e_{i}^{*})K_{j}^{-1}+(-1)^{\widehat{i}\widehat{j}}K_{i}^{-1}L(T_{1}^{-\gamma(j,i)})L(e_{j}^{*})$
$\phantom{=}\ +(-1)^{\widehat{i}(\widehat{j}+\widehat{i})}L(e_{i})L(T_{1}T_{2}^{-\gamma(j,i)})L(e_{i}^{*})L(e_{j}^{*}))(e_{i_{d-1}}\cdots e_{i_{1}})$
$=((-1)^{\widehat{i}\widehat{j}}(L(T_{1}^{-\gamma(j,i)})+(-1)^{\widehat{i}\widehat{j}}\delta_{ij}q_{i}^{-1})K_{i}^{-1}L(e_{j}^{*})$
$\phantom{=}\ +(-1)^{\widehat{i}(\widehat{j}+\widehat{i})}L(e_{i})L(T_{1}T_{2}^{-\gamma(j,i)})L(e_{i}^{*})L(e_{j}^{*}))(e_{i_{d-1}}\cdots e_{i_{1}})$
$=((-1)^{\widehat{i}\widehat{j}}(L(T_{1}^{\gamma(i,j)})+\delta_{ij}(-1)^{\widehat{i}}(q_{i}-q_{i}^{-1})+\delta_{ij}(-1)^{\widehat{i}}q_{i}^{-1})K_{i}^{-1}L(e_{j}^{*})$
$\phantom{=}\ +(-1)^{\widehat{i}(\widehat{j}+\widehat{i})}L(e_{i})L(T_{1}T_{2}^{-\gamma(j,i)})L(e_{i}^{*})L(e_{j}^{*}))(e_{i_{d-1}}\cdots e_{i_{1}})$
$=((-1)^{\widehat{i}\widehat{j}}(L(T_{1}^{\gamma(i,j)})+\delta_{ij}(-1)^{\widehat{i}}q_{i})K_{i}^{-1}L(e_{j}^{*})$
$\phantom{=}\ +(-1)^{\widehat{i}(\widehat{j}+\widehat{i})}L(e_{i})L(T_{1}T_{2}^{-\gamma(j,i)})L(e_{i}^{*})L(e_{j}^{*}))(e_{i_{d-1}}\cdots e_{i_{1}})$
$=((-1)^{\widehat{i}\widehat{j}}(L(T_{1}^{\gamma(i,j)})K_{i}^{-1}L(e_{j}^{*}))+\delta_{ij}q_{i}K_{j}^{-1}L(e_{i}^{*})$
$\phantom{=}\ +(-1)^{\widehat{i}}q_{i}^{\delta_{ij}}L(e_{i})L(T_{1}T_{2}^{-\gamma(j,i)}T_{1}^{\gamma(i,j)})L(e_{j}^{*})L(e_{i}^{*}))(e_{i_{d-1}}\cdots e_{i_{1}}),$
and
$(-1)^{\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(T_{1}^{\gamma(i,j)})L(e_{j}^{*})L(e_{i}^{*})L(e_{i})$
$=(-1)^{\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(T_{1}^{\gamma(i,j)})L(e_{j}^{*})K_{i}^{-1}+(-1)^{\widehat{i}(\widehat{j}+\widehat{i})}q_{i}^{\delta_{ij}}L(T_{1}^{\gamma(i,j)})L(e_{j}^{*})L(e_{i})L(T_{1})L(e_{i}^{*})$
$=(-1)^{\widehat{i}\widehat{j}}L(T_{1}^{\gamma(i,j)})K_{i}^{-1}L(e_{j}^{*})+(-1)^{\widehat{i}(\widehat{j}+\widehat{i})}q_{i}^{\delta_{ij}}L(T_{1}^{\gamma(i,j)})(\delta_{ij}K_{j}^{-1}$
$\phantom{=}\ +(-1)^{\widehat{i}\widehat{j}}L(e_{i})L(T_{1}^{-\gamma(j,i)})L(e_{j}^{*}))L(T_{1})L(e_{i}^{*})$
$=(-1)^{\widehat{i}\widehat{j}}L(T_{1}^{\gamma(i,j)})K_{i}^{-1}L(e_{j}^{*})+\delta_{ij}q_{i}^{\delta_{ij}}K_{j}^{-1}L(e_{i}^{*})$
$\phantom{=}\ +(-1)^{\widehat{i}}q_{i}^{\delta_{ij}}L(e_{i})L(T_{2}^{\gamma(i,j)}T_{1}^{-\gamma(j,i)}T_{2})L(e_{j}^{*})L(e_{i}^{*})$
$=(-1)^{\widehat{i}\widehat{j}}L(T_{1}^{\gamma(i,j)})K_{i}^{-1}L(e_{j}^{*})+\delta_{ij}q_{i}K_{j}^{-1}L(e_{i}^{*})$
[…]
If $i_{d}=j$, we can get the result similarly. ∎
6.2. The representations of $U_{m|n}$ on $\widehat{T}(\mathbb{V})$
For $1\leq i<m+n$ and $d\in\mathbb{N}$, we denote
$A_{i}^{(d)}=L(e_{i})L(e_{i+1}^{*})|_{\mathbb{V}^{\otimes d}},$
and
$B_{i}^{(d)}=L(e_{i+1})L(e_{i}^{*})|_{\mathbb{V}^{\otimes d}}.$
Let $\sigma:\mathbb{V}\to\mathbb{V}$ be the linear endomorphism defined by $\sigma(e_{k})=(-1)^{\widehat{k}}e_{k}$ for $1\leq k\leq m+n$.
Lemma 6.4.
We have
$A_{i}^{(d)}=A_{i}^{(1)}\otimes K_{i+1}^{-1}+\sigma^{\delta_{im}}K_{i}^{-1}\otimes A_{i}^{(d-1)},$
and
$B_{i}^{(d)}=B_{i}^{(1)}\otimes K_{i}+\sigma^{\delta_{im}}K_{i+1}\otimes B_{i}^{(d-1)}.$
Proof.
For $I=(i_{d-1},\ldots,i_{2},i_{1})\in[m+n]^{d-1},k\in[m+n]$, by direct calculation, we have
$A_{i}^{(d)}(e_{k}e_{i_{d-1}}\cdots e_{i_{1}})$
$=L(e_{i})L(e_{i+1}^{*})(e_{k}e_{i_{d-1}}\cdots e_{i_{1}})$
$=L(e_{i})L(e_{i+1}^{*})L(e_{k})(e_{i_{d-1}}\cdots e_{i_{1}})$
$=L(e_{i})(\delta_{i+1,k}K_{i+1}^{-1}+(-1)^{\widehat{i+1}\widehat{k}}L(e_{k})L(T_{1}^{-\gamma(i+1,k)})L(e_{i+1}^{*}))(e_{i_{d-1}}\cdots e_{i_{1}})$
$=L(\delta_{i+1,k}e_{i})\otimes K_{i+1}^{-1}(e_{i_{d-1}}\cdots e_{i_{1}})+(-1)^{\widehat{i+1}\widehat{k}}L(e_{i})L(e_{k})L(T_{1}^{-\gamma(i+1,k)})L(e_{i+1}^{*})(e_{i_{d-1}}\cdots e_{i_{1}})$
$=A_{i}^{(1)}(e_{k})\otimes K_{i+1}^{-1}(e_{i_{d-1}}\cdots e_{i_{1}})+(-1)^{\widehat{i+1}\widehat{k}}L(e_{i})L(e_{k})L(T_{1}^{-\gamma(i,k)-2\delta_{i,k}})L(e_{i+1}^{*})(e_{i_{d-1}}\cdots e_{i_{1}})$
$=(A_{i}^{(1)}\otimes K_{i+1}^{-1})(e_{k}e_{i_{d-1}}\cdots e_{i_{1}})+(-1)^{\widehat{i+1}\widehat{k}}(-1)^{\widehat{i}\widehat{k}}q_{i}^{\delta_{i,k}}L(e_{k}e_{i}(T_{1}^{-2})^{\delta_{i,k}})L(e_{i+1}^{*})(e_{i_{d-1}}\cdots e_{i_{1}})$
$=(A_{i}^{(1)}\otimes K_{i+1}^{-1})(e_{k}e_{i_{d-1}}\cdots e_{i_{1}})+(-1)^{\widehat{k}\delta_{im}}q_{i}^{-\delta_{i,k}}L(e_{k})L(e_{i})L(e_{i+1}^{*})(e_{i_{d-1}}\cdots e_{i_{1}})$
$=(A_{i}^{(1)}\otimes K_{i+1}^{-1})(e_{k}e_{i_{d-1}}\cdots e_{i_{1}})+(\sigma^{\delta_{i,m}}K_{i}^{-1})(e_{k})\otimes A_{i}^{(d-1)}(e_{i_{d-1}}\cdots e_{i_{1}})$
$=(A_{i}^{(1)}\otimes K_{i+1}^{-1}+\sigma^{\delta_{i,m}}K_{i}^{-1}\otimes A_{i}^{(d-1)})(e_{k}e_{i_{d-1}}\cdots e_{i_{1}}).$
The formula for $B_{i}^{(d)}$ can be proved in a similar way. ∎
Lemma 6.5.
For $1\leq i<m+n$, we have
$E_{i}K_{i}^{-1}=L(e_{i})L(e_{i+1}^{*})|_{\mathbb{V}^{\otimes d}},$
and
$K_{i}F_{i}=L(e_{i+1})L(e_{i}^{*})|_{\mathbb{V}^{\otimes d}}.$
Proof.
We only prove the first equation. The second one can be proved similarly. We use induction on $d$. It is clear for $d=1$. Suppose the equation holds for $d=r-1$, then
$E_{i}K_{i}^{-1}(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
$=(\Delta^{(r-1)}(E_{i})\Delta^{(r-1)}(K_{i}^{-1}))(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
$=((1\otimes\Delta^{(r-2)}(E_{i})+E_{i}\otimes(K_{i}K_{i+1}^{-1})^{\otimes r})((K_{i}^{-1})^{\otimes r+1}))(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
$=(K_{i}^{-1}\otimes\Delta^{(r-2)}(E_{i}K_{i}^{-1})+E_{i}K_{i}^{-1}\otimes(K_{i+1}^{-1})^{\otimes r})(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
$=((-1)^{|E_{i}|\widehat{i_{r}}}K_{i}^{-1}\otimes A_{i}^{(r-1)}+A_{i}^{(1)}\otimes K_{i+1}^{-1})(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
$=(A_{i}^{(1)}\otimes K_{i+1}^{-1}+(-1)^{\delta_{im}\widehat{i_{r}}}K_{i}^{-1}\otimes A_{i}^{(r-1)})(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
$=A_{i}^{(r)}(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})=L(e_{i})L(e_{i+1}^{*})(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}}),$
where the forth equality follows from the inductive hypothesis, and the sign comes from the super-action of $E_{i}K_{i}^{-1}$.
We further denote $E_{i,i+1}=E_{i}$, $E_{i+1,i}=F_{i}$ and
$E_{ij}=E_{ik}E_{kj}-q_{k}E_{kj}E_{ik},\quad E_{ji}=E_{jk}E_{ki}-q_{k}^{-1}E_{ki}E_{jk},\quad(i<k<j).$
Theorem 6.6.
For any $1\leq i<j\leq m+n$, we have
$E_{ij}K_{i}^{-1}=L(e_{i})L(e_{j}^{*})|_{\mathbb{V}^{\otimes d}},$
and
$K_{j}E_{ij}=L(e_{j})L(e_{i}^{*})|_{\mathbb{V}^{\otimes d}}.$
Proof.
As above, we only prove the first equation. We use induction on $j-i$. It has been showed in Lemma 6.5 for the case $j-i=1$. On $\mathbb{V}^{\otimes d}$, suppose $E_{ik}K_{i}^{-1}=L(e_{i})L(e_{k}^{*}),i<k<j$. Then
$E_{ij}K_{i}^{-1}=E_{ij}K_{i}^{-1}K_{j-1}^{-1}K_{j-1}$
$=(E_{i(j-1)}E_{(j-1)j}-q_{j-1}E_{(j-1)j}E_{i(j-1)})(K_{i}^{-1}K_{j-1}^{-1})K_{j-1}$
$=(E_{i(j-1)}K_{i}^{-1}E_{(j-1)j}K_{j-1}^{-1}-q_{j-1}E_{(j-1)j}E_{i(j-1)}K_{i}^{-1}K_{j-1}^{-1})K_{j-1}$
$=(L(e_{i})L(e_{j-1}^{*})L(e_{j-1})L(e_{j}^{*})-q_{j-1}E_{(j-1)j}L(e_{i})L(e_{j-1}^{*})K_{j-1}^{-1})K_{j-1}$
$=(L(e_{i})L(e_{j-1}^{*})L(e_{j-1})L(e_{j}^{*})-E_{(j-1)j}K_{j-1}^{-1})L(e_{i})L(e_{j-1}^{*})K_{j-1}.$
Moreover,
$(E_{(j-1)j}K_{j-1}^{-1})L(e_{i})L(e_{j-1}^{*})$
$=L(e_{j-1})L(e_{j}^{*})L(e_{i})L(e_{j-1}^{*})$
$=(-1)^{\widehat{i}\widehat{j}}L(e_{j-1})L(e_{i})L(T_{1}^{-1})L(e_{j}^{*})L(e_{j-1}^{*})$
$=(-1)^{\widehat{i}\widehat{j}+\widehat{i}\widehat{j-1}+\widehat{j}\widehat{j-1}}L(e_{i})L(e_{j-1})L(T_{1})L(e_{j-1}^{*})L(e_{j}^{*})$
$=(-1)^{\widehat{i}\widehat{j}+\widehat{i}\widehat{j-1}+\widehat{j}\widehat{j-1}+\widehat{j-1}\widehat{j-1}}L(e_{i})(L(e_{j-1}^{*})L(e_{j-1})-K_{j-1}^{-1})L(e_{j}^{*})$
$=(-1)^{(\widehat{i}+\widehat{j-1})(\widehat{j-1}+\widehat{j})}(L(e_{i})L(e_{j-1}^{*})L(e_{j-1})L(e_{j}^{*})-L(e_{i})L(e_{j}^{*})K_{j-1}^{-1})$
$=L(e_{i})L(e_{j-1}^{*})L(e_{j-1})L(e_{j}^{*})-L(e_{i})L(e_{j}^{*})K_{j-1}^{-1}.$
So we get the result. ∎
Remark 6.7.
Based on these identities on $\mathbb{V}^{\otimes d}$ and the fact that $L(e),L(e^{*}),K,K^{-1}$ all commute with $\mathscr{H}_{\infty}$, we can actually have a well-defined $\mathbf{U}_{m|n}$-module structure on $\widetilde{T}(\mathbb{V})$ and get the same result.
Furthermore, we can get the following proposition.
Proposition 6.8.
For any $v_{1},...,v_{k}\in\mathbb{V}$, $v_{1}^{*},...,v_{k}^{*}\in\mathbb{V}^{*}$, we have
$L(v_{k})\cdots L(v_{1})L(v_{1}^{*})\cdots L(v_{k}^{*})\in\rho(\mathbf{U}_{m|n}),$
where $\rho$ denotes the representation of $U_{m|n}$ on $\widetilde{T}(\mathbb{V})$.
Proof.
We first claim that
$L(e_{i})L(e_{j})L(e_{k}^{*})=aL(e_{i})L(e_{k}^{*})L(e_{j})+bL(e_{j})L(e_{k}^{*})L(e_{i})+cK_{k}^{-1}L(e_{i})+dK_{k}^{-1}L(e_{j}),$
where $a,b,c,d\in\mathbb{C}(q)$. Recall from Proposition 6.1 and 6.3 that
$L(e_{i})L(e_{j})L(T^{-\gamma(k,j)})L(e_{k}^{*})$ $=(-1)^{\widehat{j}\widehat{k}}(L(e_{i})L(e_{k}^{*})L(e_{j})-\delta_{jk}L(e_{i})K_{j}^{-1})$
$=(-1)^{\widehat{j}\widehat{k}}L(e_{i})L(e_{k}^{*})L(e_{j})+\delta_{jk}q_{k}^{-\delta_{i,k}}K_{k}^{-1}L(e_{i}).$
Then, if $i=j$, we have
$L(e_{i})L(e_{j})L(e_{k}^{*})$ $=(-1)^{\widehat{i}}q_{i}^{\gamma(k,i)}L(e_{i})L(e_{i})(L(T^{-\gamma(k,i)}))L(e_{k}^{*})$
$=q_{i}^{\gamma(k,i)}L(e_{i})(L(e_{k}^{*})L(e_{i})-\delta_{ik}K_{i}^{-1})$
$=q_{i}^{\gamma(k,i)}L(e_{i})L(e_{k}^{*})L(e_{i})-\delta_{ik}K_{k}^{-1}L(e_{i}).$
If $i\neq j$, we have
$L(e_{i})L(e_{j})L(e_{k}^{*})$
$=(-1)^{\widehat{i}\widehat{j}}L(e_{j})L(e_{i})L(T^{\gamma(i,j)})L(e_{k}^{*})$
$=(-1)^{\widehat{i}\widehat{j}}L(e_{j})L(e_{i})L(T^{-\gamma(k,i)}+c_{1})L(e_{k}^{*})$
$=(-1)^{\widehat{i}\widehat{j}}L(e_{j})L(e_{i})L(T^{-\gamma(k,i)})L(e_{k}^{*})+c_{1}L(e_{i})L(e_{j})L(T^{\gamma(j,i)})L(e_{k}^{*})$
$=(-1)^{\widehat{i}\widehat{j}}L(e_{j})L(e_{i})L(T^{-\gamma(k,i)})L(e_{k}^{*})+c_{1}L(e_{i})L(e_{j})L(T^{-\gamma(k,j)}+c_{2})L(e_{k}^{*})$
$=(-1)^{\widehat{i}\widehat{j}}L(e_{j})L(e_{i})L(T^{-\gamma(k,i)})L(e_{k}^{*})+c_{1}L(e_{i})L(e_{j})L(T^{-\gamma(k,j)})L(e_{k}^{*})+c_{1}c_{2}L(e_{i})L(e_{j})L(e_{k}^{*}),$
where $c_{1},c_{2}$ is ±$\pm(q-q^{-1})$ or 00. By direct calculation, we have $c_{1}=0$ or $c_{2}=0$ (i.e., $c_{1}c_{2}=0$), since $\gamma(i,j)=-\gamma(k,i)$ or $\gamma(j,i)=-\gamma(k,j)$. Then we get the claim. From the claim, we can see that
$L(e_{i_{k}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{*})\cdots L(e_{i_{k}}^{*})$
$=aL(e_{i_{k}})L(e_{i_{k}}^{*})\cdots\ +bL(e_{i_{k-1}})L(e_{i_{k}}^{*})\cdots\ +cK_{k}^{-1}\cdots$
where $\cdots$ are elements of the form $L(e_{i_{j-1}})\cdots L(e_{j_{1}})L(e_{j_{1}}^{*})\cdots L(e_{j_{k-1}}^{*})$ and $a,b,c\in\mathbb{C}(q)$. Continuing this procedure, we get the desired result. ∎
7. Generalized quantum Schur-Sergeev duality
7.1. Euler operators
For any nonnegative integer $k$, we denote by
$[k]=\frac{q^{k}-q^{-k}}{q-q^{-1}}\quad\mbox{and}\quad[k]!=[1][2]\cdots[k]$
the $q$-quantum integers and $q$-quantum factorials, respectively.
For $i\in[m+n],a\in\mathbb{Z}$, set
$[K_{i}:a]=\frac{q_{i}^{a}K_{i}-q_{i}^{-a}K_{i}^{-1}}{q_{i}-q_{i}^{-1}},$
$[K_{i}]^{h}_{!}=[K_{i}:0][K_{i}:-1]\cdots[K_{i}:1-h],\quad\mathrm{for}\ h\in\mathbb{N_{+}},$
and $[K_{i}]^{0}_{!}=1$. Then by Lemma 6.2 we have
(7.1) $\underbrace{L(e_{i})\cdots L(e_{i})}_{h}\underbrace{L(e_{i}^{*})\cdots L(e_{i}^{*})}_{h}=[K_{i}]^{h}_{!}.$
Set =$\mathcal{I}=\{(i_{d},i_{d-1},...,i_{1})\in[m+n]^{d}~{}|~{}i_{d}\leq i_{d-1}\leq\cdots\leq i_{1}\}$. For any $I\in\mathcal{I}$, let $[\mathbf{m}(I)]!=[\mathbf{m}_{1}(I)]![\mathbf{m}_{2}(I)]!\cdots[\mathbf{m}_{m+n}(I)]!$ where $\mathbf{m}_{k}(I)$ is the multiplicity of $k$ in $I$. By (7.1) and Proposition 6.1 we have
(7.2) $L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})=\prod_{j\in[m+n]}[K_{j}]^{\mathbf{m}_{j}(I)}_{!},\quad I=(i_{d},i_{d-1},...,i_{1})\in\mathcal{I}.$
Considering the Euler operator
$A_{d}=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*}),$
we have
Lemma 7.1.
For any $\phi\in\widetilde{T}(\mathbb{V})$,
$A_{d}(\phi)=\phi.$
Proof.
Note that, for any $I,J\in\mathcal{I}$, we have
$[K_{i}]^{\mathbf{m}_{i}(I)}_{!}.e_{J}=\begin{cases}[\mathbf{m}_{i}(I)]!e_{J},&\text{if }\mathbf{m}_{i}(I)=\mathbf{m}_{i}(J);\\[3.0pt] 0,&\text{if }\mathbf{m}_{i}(I)>\mathbf{m}_{i}(J).\end{cases}$
Then by (7.2), one has
$L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{*})\cdots L(e_{i_{d}})(e_{J})=\delta_{I,J}\Bigg{(}\prod_{i\in[m+n]}[\mathbf{m}_{i}(I)]!\Bigg{)}e_{J}=\delta_{I,J}[\mathbf{m}(I)]!e_{J}.$
Now for any $L\in[m+n]^{d}$, by (5.2), there exists $J\in\mathcal{I}$ and $T\in\mathscr{H}_{d}$ such that $e_{L}=e_{J}.T$. Since $L(e_{i})$ and $L(e_{i}^{*})$ commute with $T$, we have
$A_{d}(e_{L})=A_{d}(e_{J}).T=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}\delta_{I,J}[\mathbf{m}(I)]!e_{J}.T=e_{J}.T=e_{L}.$
7.2. Descriptions of commutants
For any $d\geq 1$ and $0\leq k\leq d$, we denote by $\mathcal{A}_{d}^{k}$ the subspace of $\mathrm{End}_{\mathbb{C}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ spanned by operators of the form
$L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{*})\cdots L(v_{d}^{*}),$
with $w_{1},\ldots,w_{d}\in\mathbb{V}$, $\sigma\in\mathscr{H}_{d+k}$, and $v_{1}^{*},\ldots,v_{k}^{*}\in\mathbb{V}^{*}$.
For any $T\in\mathscr{H}_{d+k}$, we define $\pi_{d+k}(T)\in\mathrm{End}_{\mathbb{C}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ by $\pi_{d+k}(T)(\varphi)=\varphi\cdot T$ where $\varphi\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$. Denote $\mathcal{B}_{d+k}=\pi_{d+k}(\mathscr{H}_{d+k})$.
Theorem 7.2.
Asumme that $[d]!\neq 0$. We have
$\mathrm{Comm}(\mathcal{A}_{d}^{k})=\mathcal{B}_{d+k},$
and
$\mathrm{Comm}(\mathcal{B}_{d+k})=\mathcal{A}_{d}^{k}.$
Proof.
It is easy to see that $\mathcal{A}_{d}^{k}\subset\mathrm{Comm}(\mathcal{B}_{d+k})$. Since $\mathcal{H}_{d}(q)$ is semisimple, by the dobule commutant theorem, it suffices to show $\mathrm{Comm}(\mathcal{B}_{d+k})\subset\mathcal{A}_{d}^{k}$.
For $f\in\mathrm{Comm}(\mathcal{B}_{d+k}),\phi\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have
$f(\phi)$ $=f(A_{d}(\phi))$
$=f\left(\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})\phi\right)$
$=f\left(\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}e_{i_{d}}\cdots e_{i_{1}}\tau_{I}\right)$
$=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}f(e_{i_{d}}\cdots e_{i_{1}})\tau_{I}$
$=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}L(f(e_{i_{d}}\cdots e_{i_{1}}))L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})\phi,$
where $\tau_{I}=L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})\phi\in\mathscr{H}_{d+k}$. So, $f=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}L(f(e_{i_{d}}\cdots e_{i_{1}}))L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})$. Since $L(f(e_{i_{d}}\cdots e_{i_{1}}))\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have $f\in\mathcal{A}_{d}^{k}$. ∎
Theorem 7.3 (Mitsuhashi).
Asumme that $[d]!\neq 0$. For each $d\geq 1$, we have
$\mathrm{Comm}(\pi_{d}(\mathscr{H}_{d}(q)))=\rho_{d}(\mathbf{U}_{m|n}),$
and
$\mathrm{Comm}(\pi_{d}(\mathscr{H}_{d}(q)))=\rho_{d}(\mathbf{U}_{m|n}),$
where $\rho_{d}$ denotes the representation of $\mathbf{U}_{m|n}$ on $\mathbb{V}^{\otimes d}\cong\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d}$.
Proof.
We just need to prove that $\rho_{d}(\mathbf{U}_{m|n})=\mathcal{A}_{d}^{0}$.
Proposition 6.8 shows that $\mathcal{A}_{d}^{0}\subset\rho_{d}(\mathbf{U}_{m|n})$. The other side is proved as follows. We first notice that $K_{i}\in\mathrm{Comm}(\pi_{d}(\mathscr{H}_{d}(q)))$. In the proof of Theorem 7.2, we have seent that $K_{i}\in\mathcal{A}_{d}^{0}$. Actually, we have
$K_{i}=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}q_{j}^{\delta_{ii_{d}}+\cdots+\delta_{ii_{1}}}L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*}).$
Since Theorem 6.6 shows that $E_{ij}=L(e_{i})L(e_{j}^{*})K_{i}$ and $E_{ij}=K_{j}^{-1}L(e_{j})L(e_{i}^{*})$ for $i<j$ on ${\mathbb{V}^{\otimes d}}$, these elements commute with $\mathcal{B}_{d}$, then $\rho_{d}(\mathbf{U}_{m|n})\subset\mathcal{A}_{d}^{0}$. ∎

### v2

@@
 Abstract.
-Following [It12] and [It15], we construct two super-extensions of the usual tensor algebra through the super-actions of symmetric groups and Hecke algebras respectively.
-For each extension, we consider a special type of derivations coming from covectors, and study the the space generated, in some special manner, by these derivations and operators from left multiplication by vectors and permutations.
-Duality theorems of these spaces and the super-actions are proved.
-As an application, we provide a new proof of the Schur-Sergeev duality theorem, as well as its quantum version.
+We construct a super-extension of the usual $q$-tensor algebra through super-actions of Hecke algebras.
+A double centralizer property is established on this extension space, which provides a modest generalization of the quantum Schur-Sergeev duality.
+As an application, we present a new proof of the quantum Schur-Sergeev duality.
 1. Introduction
 Denote by $V^{\otimes d}$ the $d$-fold of a vector space $V$.
-In [Sc27], Schur considered the relation between the permutation action of the symmetric group $\mathfrak{S}_{d}$ and the diagonal action of the general linear group $\mathrm{GL(V)}$ on $V^{\otimes d}$.
-This is the well-known Schur-Weyl duality nowadays.
-It plays a fundamental role in representation theory (see, e.g., [We46]).
-Since its importance, this duality theorem has been generalized in several directions during the past decades.
-For example, in the late 1980s, Sergeev [Ser85] and Berele and Regev [BR87] discovered independently a super version of the Schur-Weyl duality.
-Meanwhile, Jimbo [Jim86] introduced a duality theorem about Hecke algebras and quantum enveloping algebras.
-In [Mi06], Mitsuhashi extended Jimbo’s theorem to the super case.
+In [Sc27], Schur studied the relation between the permutation action of the symmetric group $\mathfrak{S}_{d}$ and the diagonal action of the general linear group $\mathrm{GL}(V)$ on $V^{\otimes d}$.
+He pointed out that the actions of these two groups on $V^{\otimes d}$ admit a double centralizer property, which is just the celebrated Schur-Weyl duality.
+Since its importance in representation theory, this duality has been generalized in several directions during the past decades.
+In the late 1980s, the Schur-Sergeev duality, which is a super version of the Schur-Weyl duality, was established by Sergeev [Ser85] (also by Berele and Regev [BR87] independently).
+Meanwhile, Jimbo [Jim86] introduced a duality between Hecke algebras and general linear quantum groups, which is called the Schur-Jimbo duality nowadays.
+In [Mi06], Mitsuhashi provided a super quantized version of the Schur-Weyl duality, which can be regarded as a common generalization of the dualities due to Sergeev and Jimbo.
 Usually, in order to prove these dualities, people use the double commutant theorem.
-For example, for the classical Schur-Weyl duality, one needs to show that $\mathrm{GL}_{d}(\mathbb{C})=\mathrm{End}_{\mathfrak{S}_{d}}(\mathbb{C}^{d})$.
-Then by the semisimplicity of the group algebra $\mathbb{C}\mathfrak{S}_{d}$ one has $\mathbb{C}\mathfrak{S}_{d}=\mathrm{End}_{\mathrm{GL}_{d}(\mathbb{C})}(\mathbb{C}^{d})$.
-The verification of the inclusion $\mathrm{GL}_{d}(\mathbb{C})\subset\mathrm{End}_{\mathfrak{S}_{d}}(\mathbb{C}^{d})$ is trivial.
+For example, for the classical Schur-Weyl duality, one needs to show that $\mathbb{C}\mathrm{GL}(V)=\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d}}(V^{\otimes d})$.
+Then by the semisimplicity of the group algebra $\mathbb{C}\mathfrak{S}_{d}$ one has $\mathbb{C}\mathfrak{S}_{d}=\mathrm{End}_{\mathbb{C}\mathrm{GL}(V)}(V^{\otimes d})$.
+The verification of the inclusion $\mathbb{C}\mathrm{GL}(V)\subset\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d}}(V^{\otimes d})$ is trivial.
 But the reverse direction requires further non-trivial techniques.
@@
 In [It12], Itoh generalized the usual tensor algebra $T(V)$ by considering actions of symmetric groups.
-He studied operators of left multiplication by vectors and linear functionals, and established a duality theorem between the space generated by these operators and the infinity symmetric group.
-As a consequence, he provided a new proof of the Schur-Weyl dual theorem.
-In a subsequent paper [It15], he generalized his constructions to quantum case in which symmetric group actions are replaced by Hecke algebra actions.
-Using a similar argument, he gave a new proof of Jimbo’s duality theorem.
-The main purpose of this paper is to generalize Itoh’s results to super cases.
-First of all, we consider the right super-action of $\mathfrak{S}_{d}$ on $d$-fold of a vector superspace $V=V_{\overline{0}}\oplus V_{\overline{1}}$.
-Using this action, we define an associative algebra structure on the space
-$\overline{T}(V)=\bigoplus_{d\geq 0}V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathfrak{S}_{\infty}.$
-For each covector $v^{\ast}\in V^{\ast}$, we define a derivation $L(v^{\ast})$ of $\overline{T}(V)$.
-Denote by $L(w)$ and $L(\sigma)$ the operators coming from left multiplication by a vector $w$ and a permutation $\sigma$ respectively.
-We show that the space $\mathcal{A}_{d}^{k}\subset\mathrm{End}_{\mathbb{C}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$ spanned by operators of the form
-$L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{\ast})\cdots L(v_{d}^{\ast}),$
-is the commutant of $\mathfrak{S}_{d+k}$.
-Notice that $\mathcal{A}_{d}^{0}$ is just action of the Lie superalgebra $\mathfrak{gl}(m|n)$.
-We provide a new proof of the Schur-Sergeev duality.
-Secondly, we consider the right super-action of the Hecke algebra $\mathscr{H}_{d}$ on $V^{\otimes d}$, and construct an algebra
-$\widetilde{T}(V)=\bigoplus_{d\geq 0}V^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{\infty}.$
-Similar duality in this algebra is proved. As an application, we provide a new proof of Mitsuhashi’s duality.
+He studied operators of left multiplication by vectors and linear functionals, and established a duality theorem between the algebra generated by these operators and the infinity symmetric group.
+As a consequence, he provided a new proof of the Schur-Weyl duality theorem.
+In a subsequent paper [It15], he generalized his construction to quantum case in which symmetric group actions are replaced by Hecke algebra actions.
+Using a similar argument, he gave a new proof of Schur-Jimbo duality.
+The main purpose of this paper is to generalize Itoh’s results to the quantum super case.
+First, we establish the relevant construction for the super case and obtain corresponding results.
+Then, based on the special challenges of the quantum situation, we achieve the quantum super case through a generally similar but slightly different approach.
+For the $q$-tensor superspace $\mathbb{V}^{\otimes d}$, we introduce a map $\gamma$ to simplify and unify the parity calculation problems that arise in the action of the Hecke algebra $\mathscr{H}_{d}$.
+Next, we construct the quantum super-extension algebra
+$\widetilde{T}(\mathbb{V})=\bigoplus_{d\geq 0}\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{\infty},$
+and define operators $L(\phi)$, $L(v^{\ast})$ over it.
+Following this, we carefully verify the well-definedness of the relevant operators and compute their commutation relations.
+Finally, in Theorem  7.2 we demonstrate a Schur-Weyl type duality
+$\mathcal{A}_{d}^{k}\curvearrowright\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}\curvearrowleft\mathscr{H}_{d+k},$
+where $\mathcal{A}_{d}^{k}$ is the subalgebra of $\mathrm{End}_{\mathbb{C}(q)}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ spanned by operators of the form
+$L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{\ast})\cdots L(v_{d}^{\ast})$
+with $w_{1},\ldots,w_{d}\in\mathbb{V}$, $\sigma\in\mathscr{H}_{d+k}$, and $v_{1}^{*},\ldots,v_{k}^{*}\in\mathbb{V}^{*}$.
+In the special case of $k=0$, we recover Mitsuhashi’s duality, and Schur-Sergeev duality is a corollary for the non-quantum case.
 This paper is organized as follows.
-We construct the super-extension $\overline{T}(V)$ in Section 2 and study the communication relations among $L(v^{\ast})$, $L(w)$, and $L(\sigma)$ in Section 3.
-Duality theorems are proved in Section 4.
-We deal with the the quantum version $\widetilde{T}(V)$ in Section 5, 6, and 7.
+In Sections 2 - 4, we construct the super-extension algebra, study the operators defined over it and prove the duality theorems for super case finally.
+In Sections 5 - 7, we deal with the quantum super version.
 Acknowledgement
-We would like to thank Li Luo for many helpful discussions. Especially, he introduced the Schur-Sergeev duality and its quantum version to us which initiate this work.
+We would like to thank Li Luo for many helpful discussions.
+Especially, he introduced the Schur-Sergeev duality and its quantum version to us which initiates this work.
+We would also like to thank the referee for thorough review and helpful suggestions.
 2. Super-extensions through symmetric groups
 2.1. Symmetric groups
-For any positive integer $d$, we denote by $\mathfrak{S}_{d}$ the symmetric group of the set $[d]=\{1,2,\ldots,d\}$, and by $s_{i}$, $11\leq i\leq d-$ 1, the transposition interchanging $i$ and $i+1$.
+For any positive integer $d$, we denote by $\mathfrak{S}_{d}$ the symmetric group of the set $d]=\{1,2,\ldots,d\}$, and by $s_{i}$, $1\leq i\leq d-1$, the transposition interchanging $i$ and $i+1$.
 We often use cycle notation to denote a permutation.
-So $\tau=(i_{1}\ i_{2}\ \cdots\ i_{k})\in\mathfrak{S}_{d}$ maps $i_{r}$ to $i_{r+1}$ for $11\leq r\leq k-$ 1 and $i_{r}$ to $i_{1}$, and fixes other numbers.
+So $\tau=(i_{1}\ i_{2}\ \cdots\ i_{k})\in\mathfrak{S}_{d}$ maps $i_{r}$ to $i_{r+1}$ for $1\leq r\leq k-1$ and $i_{k}$ to $i_{1}$, and fixes other numbers.
 There is a natural sequence of inclusions: $\mathfrak{S}_{1}\subset\cdots\subset\mathfrak{S}_{d}\subset\cdots$.
@@
 $s_{i}s_{j}=s_{j}s_{i},\text{ whenever }|i-j|>1.$
-Given $k\geq 1$, we define a group homomorphism :${}^{\uparrow k}:\mathfrak{S}_{\infty}\rightarrow\mathfrak{S}_{\infty}$ by $s_{i}^{\uparrow k}=s_{i+k}$.
+Given $k\geq 1$, we define a group homomorphism $^{\uparrow k}:\mathfrak{S}_{\infty}\rightarrow\mathfrak{S}_{\infty}$ by $s_{i}^{\uparrow k}=s_{i+k}$.
 2.2. Superspaces and general linear Lie superalgebras
@@
 Vectors in $V_{\overline{0}}$ (resp.
-$V_{\overline{1}}$) are of parity 00 (resp.
+$V_{\overline{1}}$) are of parity 0 (resp.
 parity $1$).
@@
 In this section, we will denote by $e_{i}$, $1\leq i\leq m+n$, the vector in $V$ that is 1 in the $i$-th position and 0 elsewhere.
-Then $\{e_{1},\ldots,e_{m}\}$ and $\{e_{m+1},\ldots,e_{m+n}\}$ are bases of $V_{\overline{0}}$ and $V_{\overline{1}}$ respectively.
-We denote by $\{e_{1}^{\ast},\ldots,e_{m+n}^{\ast}\}$ the dual basis of $\{e_{1},\ldots,e_{m+n}\}$, i.e., $e_{i}^{\ast}(e_{j})=\delta_{ij}$.
+Then $\{e_{1},\ldots,e_{m}\}$ and $\{e_{m+1},\ldots,e_{m+n}\}$ are bases of $V_{\overline{0}}$ and $V_{\overline{1}}$, respectively.
+We denote $\{e_{1}^{\ast},\ldots,e_{m+n}^{\ast}\}$ as the dual basis of $\{e_{1},\ldots,e_{m+n}\}$ in the dual space $V^{\ast}$, i.e., $e_{i}^{\ast}(e_{j})=\delta_{ij}$.
 Here $\delta_{ij}$ is the Kronecker symbol.
@@
 The general linear Lie superalgebra $\mathfrak{gl}(m|n)$ is
-$\bigg{\{}\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\bigg{|}A\in M_{m\times m}(\mathbb{C}),B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg{\}}.$
+$\bigg\{\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\bigg|~A\in M_{m\times m}(\mathbb{C}),B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg\}.$
 We can see that
@@
 where
-$\mathfrak{gl}(m|n)_{\overline{0}}=\bigg{\{}\left(\begin{array}[]{cc}A&0\\ 0&D\end{array}\right)\bigg{|}A\in M_{m\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg{\}},$
+$\mathfrak{gl}(m|n)_{\overline{0}}=\bigg\{\left(\begin{array}[]{cc}A&0\\ 0&D\end{array}\right)\bigg|~A\in M_{m\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg\},$
 and
-$\mathfrak{gl}(m|n)_{\overline{1}}=\bigg{\{}\left(\begin{array}[]{cc}0&B\\ C&0\end{array}\right)\bigg{|}B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C})\bigg{\}}.$
+$\mathfrak{gl}(m|n)_{\overline{1}}=\bigg\{\left(\begin{array}[]{cc}0&B\\ C&0\end{array}\right)\bigg|~B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C})\bigg\}.$
 The parity of a homogeneous element $g\in\mathfrak{gl}(m|n)$ is denoted by $|g|$.
-As usual, for any $1\leq i,j\leq m+n$, let $E_{ij}$ be the $(m+n)\times(m+n)$ matrix with $1$ in the $(i,j)$-position and 00 elsewhere.
+As usual, for any $1\leq i,j\leq m+n$, let $E_{ij}$ be the $(m+n)\times(m+n)$ matrix with $1$ in the $(i,j)$-position and 0 elsewhere.
 Then $|E_{ij}|=|e_{i}|+|e_{j}^{\ast}|$.
@@
 where $g\in\mathfrak{gl}(m|n)$ and $v_{1},\ldots,v_{d}\in V$.
-On the other hand, the symmetric group $\mathfrak{S}_{d}$ acts from right on $[m+n]^{d}=\{(i_{d},\ldots,i_{1})|1\leq i_{1},\ldots,i_{d}\leq m+n\}$ as follows: for each $1\leq k<d$ and $I\in[m+n]^{d}$,
-$I.s_{k}=(i_{d},\ldots,i_{k},i_{k+1},\ldots,i_{1}).$
-Thus $I.\sigma=(i_{\sigma(d)},\ldots,i_{\sigma(1)})$ for $\sigma\in\mathfrak{S}_{d}$. We then have a right $\mathfrak{S}_{d}$-action on $V^{\otimes d}$ defined by
-$(v_{d}\otimes\cdots\otimes v_{1}).s_{i}=(-1)^{|v_{i}|\cdot|v_{i+1}|}v_{d}\otimes\cdots\otimes v_{i}\otimes v_{i+1}\otimes\cdots\otimes v_{1}.$
+On the other hand, the symmetric group $\mathfrak{S}_{d}$ acts from right on $m+n]^{d}=\{(i_{d},\ldots,i_{1})|1\leq i_{1},\ldots,i_{d}\leq m+n\}$ as follows: for each $1\leq k<d$ and $I=(i_{d},\ldots,i_{1})\in[m+n]^{d}$,
+$I\cdot s_{k}=(i_{d},\ldots,i_{k},i_{k+1},\ldots,i_{1}).$
+Thus $I\cdot\sigma=(i_{\sigma(d)},\ldots,i_{\sigma(1)})$ for $\sigma\in\mathfrak{S}_{d}$. We then have a right $\mathfrak{S}_{d}$-action on $V^{\otimes d}$ defined by
+$(v_{d}\otimes\cdots\otimes v_{1})\cdot s_{i}=(-1)^{|v_{i}|\cdot|v_{i+1}|}v_{d}\otimes\cdots\otimes v_{i}\otimes v_{i+1}\otimes\cdots\otimes v_{1}.$
 For example, if $\sigma=(123)=s_{1}s_{2}$, then
-$(v_{3}\otimes v_{2}\otimes v_{1}).\sigma=(-1)^{|v_{2}|\cdot|v_{1}|+|v_{3}|\cdot|v_{1}|}v_{1}\otimes v_{3}\otimes v_{2}.$
+$(v_{3}\otimes v_{2}\otimes v_{1})\cdot\sigma=(-1)^{|v_{2}|\cdot|v_{1}|+|v_{3}|\cdot|v_{1}|}v_{1}\otimes v_{3}\otimes v_{2}.$
 In general, for any $\sigma\in\mathfrak{S}_{d}$,
-(2.1) $(v_{d}\otimes v_{d-1}\otimes\cdots\otimes v_{1}).\sigma=\varepsilon_{v_{d},\ldots,v_{1};\sigma}v_{\sigma(d)}\otimes v_{\sigma(d-1)}\otimes\cdots\otimes v_{\sigma(1)},$
-where $\varepsilon_{v_{d},\ldots,v_{1};\sigma}$ is either $1$ or $-1$, depending on $v_{d},\ldots,v_{1}$ and $\sigma$.
+(2.1) $(v_{d}\otimes v_{d-1}\otimes\cdots\otimes v_{1})\cdot\sigma=\varepsilon_{v_{d},\ldots,v_{1};\sigma}v_{\sigma(d)}\otimes v_{\sigma(d-1)}\otimes\cdots\otimes v_{\sigma(1)},$
+where $\varepsilon_{v_{d},\ldots,v_{1};\sigma}$ is either $1$ or 1-1, depending on $v_{d},\ldots,v_{1}$ and $\sigma$.
 2.4. Super-extensions
@@
 $\overline{T}(V)=\bigoplus_{d\geq 0}\overline{T}_{d}(V).$
-We define an associative product $\cdot$ on $\overline{T}$ as follows: for any $u_{k}\cdots u_{1}\otimes\tau\in\widehat{T}_{k}(V)$ and $v_{d}\cdots v_{1}\otimes\sigma\in\overline{T}_{d}(V)$,
-$(u_{k}\cdots u_{1}\otimes\tau)\cdot(v_{d}\cdots v_{1}\otimes\sigma)=u_{k}\cdots u_{1}v_{d}\cdots\otimes v_{1}\otimes\tau^{\uparrow d}\sigma.$
+We define an associative product $\cdot$ on $\overline{T}(V)$ as follows: for any $u_{k}\cdots u_{1}\otimes\tau\in\overline{T}_{k}(V)$ and $v_{d}\cdots v_{1}\otimes\sigma\in\overline{T}_{d}(V)$,
+$(u_{k}\cdots u_{1}\otimes\tau)\cdot(v_{d}\cdots v_{1}\otimes\sigma)=u_{k}\cdots u_{1}v_{d}\cdots v_{1}\otimes\tau^{\uparrow d}\sigma.$
 For any $v,w\in V$, we have immediately that
@@
 For any $\varphi\in\overline{T}(V)$, we define
-$\begin{array}[]{cccc}L(\varphi):&\overline{T}(V)&\rightarrow&\overline{T},\\[3.0pt] &\psi&\mapsto&\varphi\cdot\psi.\end{array}$
+$\begin{array}[]{cccc}L(\varphi):&\overline{T}(V)&\rightarrow&\overline{T}(V),\\[3.0pt] &\psi&\mapsto&\varphi\cdot\psi.\end{array}$
 For a homogeneous element $v^{\ast}\in V^{\ast}$, we define $L(v^{\ast}):\overline{T}(V)\rightarrow\overline{T}(V)$ by
 $L(v^{\ast})(v_{d}\cdots v_{1}\otimes\sigma)$
-$=\sum_{k=1}^{d}(-1)^{|v^{\ast}|(|v_{d}|+\cdots+|v_{k+1}|)}\langle v^{\ast},v_{k}\rangle v_{d}\cdots\hat{v}_{k}\cdots v_{1}\otimes(d\ d-1\ \cdots\ k+1\ k)\sigma,$
-$=\sum_{k=1}^{d}\Big{(}(v^{\ast}\otimes\mathrm{id}_{V^{\otimes d-1}})\big{(}(v_{d}\cdots v_{1}).(s_{k}\cdots s_{d-2}s_{d-1})\big{)}\Big{)}\otimes s_{d-1}s_{d-2}\cdots s_{k}\sigma$
+$=\sum_{k=1}^{d}(-1)^{|v^{\ast}|(|v_{d}|+\cdots+|v_{k+1}|)}\langle v^{\ast},v_{k}\rangle v_{d}\cdots\hat{v}_{k}\cdots v_{1}\otimes(d\ d-1\ \cdots\ k+1\ k)\sigma$
+$=\sum_{k=1}^{d}\Big((v^{\ast}\otimes\mathrm{id}_{V^{\otimes d-1}})\big((v_{d}\cdots v_{1})\cdot(s_{k}\cdots s_{d-2}s_{d-1})\big)\Big)\otimes s_{d-1}s_{d-2}\cdots s_{k}\sigma,$
 where the symbol $\hat{v}_{k}$ means the omission of $v_{k}$. The well-definedness of these two maps can be checked directly and stated formally as:
 Proposition 2.1.
-The operators $L(\sigma)$, $L(e_{i})$, and $L(e_{i}^{\ast})$ for $\sigma\in\mathfrak{S}_{\infty},1\leq i\leq m+n$ are elements in $\mathrm{End}_{\mathfrak{S}_{\infty}}(\overline{T}(V))$, where $\mathfrak{S}_{\infty}$ acts on $\overline{T}(V)$ by right multiplication.
+The operators $L(\sigma)$, $L(e_{i})$, and $L(e_{i}^{\ast})$ for $\sigma\in\mathfrak{S}_{\infty},1\leq i\leq m+n$ are elements in $\mathrm{End}_{\mathbb{C}\mathfrak{S}_{\infty}}(\overline{T}(V))$, where $\mathfrak{S}_{\infty}$ acts on $\overline{T}(V)$ by right multiplication.
 3. Properties of operators
@@
 By direct calculation, we have
-$L(e_{i})L(e_{j}^{\ast})(v_{d}\otimes\cdots\otimes v_{1}\otimes\sigma)$
+$L(e_{i})L(e_{j}^{\ast})(v_{d}\cdots v_{1}\otimes\sigma)$
 $=\sum_{k=1}^{d}(-1)^{|e_{j}^{\ast}|(|v_{d}|+\cdots+|v_{k+1}|)}\langle e_{j}^{\ast},v_{k}\rangle e_{i}v_{d}\cdots\hat{v}_{k}\cdots v_{1}\otimes s_{d-1}s_{d-2}\cdots s_{k}\sigma$
@@
 Set
-$\mathcal{S}_{1}$ =$=\{1,s_{1}\},$
-$\mathcal{S}_{2}$ =$=\{1,s_{2},s_{2}s_{1}\},$
-$\mathcal{S}_{3}$ =$=\{1,s_{3},s_{3}s_{2},s_{3}s_{2}s_{1}\}$
+$\mathcal{S}_{1}$ $=\{1,s_{1}\},$
+$\mathcal{S}_{2}$ $=\{1,s_{2},s_{2}s_{1}\},$
+$\mathcal{S}_{3}$ $=\{1,s_{3},s_{3}s_{2},s_{3}s_{2}s_{1}\},$
 $\vdots$
-$\mathcal{S}_{d-1}$ =$=\{1,s_{d-1},s_{d-1}s_{d-2},\ldots,s_{d-1}s_{d-2}\cdots s_{2}s_{1}\}.$
+$\mathcal{S}_{d-1}$ $=\{1,s_{d-1},s_{d-1}s_{d-2},\ldots,s_{d-1}s_{d-2}\cdots s_{2}s_{1}\}.$
 Then every $\sigma\in\mathfrak{S}_{d}$ can be written uniquely as
@@
 $\ \ \ \ \ \ \ \ \ \ \ \ (e_{i_{\tau(d)}}^{\ast}\otimes e_{i_{\tau(d-1)}}^{\ast}\otimes\mathrm{id}_{V^{\otimes d-2}})\Bigg(\sum_{\sigma_{2}\in\mathcal{S}_{d-2}}\big((e_{i_{d}}\cdots e_{i_{1}}).\sigma_{d-1}^{-1}\big).\sigma_{d-2}^{-1}\otimes\sigma_{d-2}\sigma_{d-1}\Bigg)$
-=$\cdots$= ⋯
+$=\cdots$
 $=\sum_{\sigma_{1}\in\mathcal{S}_{1}}\cdots\sum_{\sigma_{d-1}\in\mathcal{S}_{d-1}}(e_{i_{\tau(d)}}^{\ast}\cdots e_{i_{\tau(1)}}^{\ast})\big((e_{i_{d}}\cdots e_{i_{1}}).(\sigma_{d-1}^{-1}\cdots\sigma_{1}^{-1})\big)\otimes\sigma_{1}\cdots\sigma_{d-1}$
@@
 $=|(\mathfrak{S}_{d})_{I}|e_{i_{d}}\cdots e_{i_{1}}$
-$=\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}}.$
+$=\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}},$
+where the fifth equality follows by an isomorphism $(\mathfrak{S}_{d})_{I.\tau}\to(\mathfrak{S}_{d})_{I},\sigma\mapsto\tau\sigma\tau^{-1}$. ∎
 Corollary 3.4.
 For any $I=(i_{d},\ldots,i_{1}),J=(j_{d},\ldots,j_{1})\in[m+n]^{d}$, we have
-$L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(e_{j_{d}}\cdots e_{j_{1}})=\begin{cases}\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}},&\text{if }\mathbf{m}(I)=\mathbf{m}(J),\\[3.0pt] 0,&\text{if }\mathbf{m}(I)\neq\mathbf{m}(J).\end{cases}$
+$L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(e_{j_{d}}\cdots e_{j_{1}})=\begin{cases}\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}}&\text{if }\mathbf{m}(I)=\mathbf{m}(J),\\[3.0pt] 0&\text{if }\mathbf{m}(I)\neq\mathbf{m}(J).\end{cases}$
 We provide two more interesting formulas of such type which will be used in the sequel.
@@
 $=\frac{1}{d!}\sum_{(i_{d},\ldots,i_{3})\in[m+n]^{d-2}}L(e_{i_{d}})\cdots L(e_{i_{3}})\big(2L(e_{i_{3}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\big)$
-=$\cdots$= ⋯
+$=\cdots$
 $=\frac{1}{d!}d!\varphi$
@@
 4. Generalizations of the Schur-Sergeev duality
-Let $W$ be a vector space and $A$ a subset of $\mathrm{End}_{\mathbb{C}}(W)$. The commutant of $A$ is denoted by $\mathrm{Comm}(A)$. That is
-$\mathrm{Comm}(A)=\{f\in\mathrm{End}_{\mathbb{C}}(W)|fg=gf\text{ for all }g\in A\}.$
+Let $W$ be a vector space and $A$ a subset of $\mathrm{End}_{\mathbb{C}}(W)$. The commutant of $A$ is denoted by $\mathrm{End}_{A}(W)$. That is
+$\mathrm{End}_{A}(W)=\{f\in\mathrm{End}_{\mathbb{C}}(W)~|~fg=gf\text{ for all }g\in A\}.$
 For any $d\geq 1$ and $0\leq k\leq d$, we consider $\mathrm{End}_{\mathbb{C}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$.
@@
 with $w_{1},\ldots,w_{d}\in V$, $\sigma\in\mathfrak{S}_{d+k}$, and $v_{1}^{\ast},\ldots,v_{k}^{\ast}\in V^{\ast}$.
-For any $t\in\mathbb{C}\mathfrak{S}_{d+k}$, we define $\pi_{d+k}(t)\in\mathrm{End}_{\mathbb{C}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$ by $\pi_{d+k}(t)(\varphi)=\varphi\cdot t$ where $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$.
-Denote $\mathcal{B}_{d+k}=\pi_{d+k}(\mathbb{C}\mathfrak{S}_{d+k})$.
+The representation $\pi_{d+k}$ of $\mathbb{C}\mathfrak{S}_{d+k}$ on $V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$ is given by right multiplication and we identify $\pi_{d+k}(\mathbb{C}\mathfrak{S}_{d+k})$ with $\mathbb{C}\mathfrak{S}_{d+k}$.
 Theorem 4.1.
-Keep the notations above. We have
-$\mathrm{Comm}(\mathcal{A}_{d}^{k})=\mathcal{B}_{d+k},$
+We have
+$\mathcal{A}_{d}^{k}=\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d+k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$
 and
-$\mathrm{Comm}(\mathcal{B}_{d+k})=\mathcal{A}_{d}^{k}.$
+$\mathrm{End}_{\mathcal{A}_{d}^{k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})=\mathbb{C}\mathfrak{S}_{d+k}.$
 Proof.
-Apparently, by Proposition 2.1, we have $\mathcal{A}_{d}^{k}\subset\mathrm{Comm}(\mathcal{B}_{d+k})$.
-Now for any $f\in\mathrm{Comm}(\mathcal{B}_{d+k})$ and $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$, it follows from Proposition 3.6 that
-$f(\varphi)$
-$=f(A_{d}(\varphi))$
+Apparently, by Proposition 2.1, we have $\mathcal{A}_{d}^{k}\subset\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d+k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$.
+Now for any $f\in\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d+k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$ and $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$, it follows from Proposition 3.6 that
+$f(\varphi)$ $=f(A_{d}(\varphi))$
 $=f\Bigg(\frac{1}{d!}\sum_{(i_{d},\ldots,i_{1})\in[m+n]^{d}}L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi)\Bigg)$
@@
 $=\sum_{J=(j_{d},\ldots,j_{1})\in[m+n]^{d}}L(e_{j_{d}})\cdots L(e_{j_{1}})L(t_{J})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(\varphi).$
-Hence $f\in\mathcal{A}_{d}^{k}$. As a consequence, $\mathrm{Comm}(\mathcal{B}_{d+k})\subset\mathcal{A}_{d}^{k}$.
-Since $\mathbb{C}\mathfrak{S}_{d+k}$ is a semisimple algebra, it follows from the double commutant theorem that $\mathrm{Comm}(\mathcal{A}_{d}^{k})=\mathcal{B}_{d+k}$.∎
-As an application of the above theorem, we provide a new proof of the Schur-Sergeev duality theorem. Firstly, we need the following technical lemma.
+Hence $f\in\mathcal{A}_{d}^{k}$.
+As a consequence, $\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d+k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})\subset\mathcal{A}_{d}^{k}$.
+Since $\mathbb{C}\mathfrak{S}_{d+k}$ is a semisimple algebra, it follows from the double commutant theorem that $\mathrm{End}_{\mathcal{A}_{d}^{k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})=\mathbb{C}\mathfrak{S}_{d+k}$.∎
+As an application of the above theorem, we provide a new proof of the Schur-Sergeev duality. Firstly, we need the following technical lemma.
 Lemma 4.2.
@@
 $L(e_{i})L(e_{j})L(e_{k}^{\ast})=(-1)^{|e_{i}|(|e_{j}|+|e_{k}^{\ast}|)}\big(L(e_{j})L(e_{k}^{\ast})L(e_{i})-\delta_{ik}L(e_{j})\big),$
-on $\overline{T}_{d}$.
+on $\overline{T}_{d}(V)$.
 Proof.
@@
 as desired.∎
-Theorem 4.3 (Schur-Sergeev duality).
+Corollary 4.3 (Schur-Sergeev duality).
 For each $d\geq 1$, we have
-$\mathrm{Comm}\big{(}\rho_{d}(U(gl(m|n)))\big{)}=\mathcal{B}_{d},$
+$\rho_{d}(U(\mathfrak{gl}(m|n)))=\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d}}(V^{\otimes d})$
 and
-$\mathrm{Comm}(\mathcal{B}_{d})=\rho_{d}(U(gl(m|n))).$
+$\mathrm{End}_{\rho_{d}(U(\mathfrak{gl}(m|n)))}(V^{\otimes d})=\mathbb{C}\mathfrak{S}_{d},$
+where $\rho_{d}$ denotes the representation of $U(\mathfrak{gl}(m|n))$ on $V^{\otimes d}\cong V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d}$.
 Proof.
-We are going to show that $\mathcal{A}_{d}^{0}=\rho_{d}(U(gl(m|n)))$. We first notice that every element in $\mathcal{A}_{d}^{0}$ is a linear combination of operators of the form
+We are going to show that $\mathcal{A}_{d}^{0}=\rho_{d}(U(\mathfrak{gl}(m|n)))$. We first notice that every element in $\mathcal{A}_{d}^{0}$ is a linear combination of operators of the form
 $L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{\ast})\cdots L(v_{d}^{\ast}),$
@@
 $=\frac{1}{(d-1)!}\sum_{k_{1},\ldots,k_{d-1}=1}^{m+n}L(e_{i})L(e_{k_{d-1}})\cdots L(e_{k_{1}})L(e_{k_{1}}^{\ast})\cdots L(e_{k_{d-1}}^{\ast})L(e_{j}^{\ast})(\varphi).$
-Therefore $\rho_{d}(E_{ij})\in\mathcal{A}_{d}^{0}$, and hence $\rho_{d}(U(gl(m|n)))\subset\mathcal{A}_{d}^{0}$.
-On the other hand, for any $(i_{p},\ldots,i_{1}),(j_{p},\ldots,j_{1})\in[m+n]^{p}$, we use induction on $p$ to show that $L(e_{i_{p}})\cdots L(e_{i_{1}})L(e_{j_{1}}^{\ast})\cdots L(e_{j_{p}}^{\ast})\in\rho_{d}(U(gl(m|n)))$.
+Therefore $\rho_{d}(E_{ij})\in\mathcal{A}_{d}^{0}$, and hence $\rho_{d}(U(\mathfrak{gl}(m|n)))\subset\mathcal{A}_{d}^{0}$.
+On the other hand, for any $(i_{p},\ldots,i_{1}),(j_{p},\ldots,j_{1})\in[m+n]^{p}$, we use induction on $p$ to show that $L(e_{i_{p}})\cdots L(e_{i_{1}})L(e_{j_{1}}^{\ast})\cdots L(e_{j_{p}}^{\ast})\in\rho_{d}(U(\mathfrak{gl}(m|n)))$.
 The case $p=1$ follows from (3.1).
@@
 Let $\mathscr{H}_{\infty}$ be the inductive limit of the natural inclusions $\mathscr{H}_{0}\subset\mathscr{H}_{1}\subset\cdots$.
-Given $k\geq 1$, we define an algebra homomorphism :${}^{\uparrow k}:\mathscr{H}_{\infty}\rightarrow\mathscr{H}_{\infty}$ by $T_{i}^{\uparrow k}=T_{i+k}$.
+Given $k\geq 1$, we define an algebra homomorphism $^{\uparrow k}$$\mathscr{H}_{\infty}\rightarrow\mathscr{H}_{\infty}$ by $T_{i}^{\uparrow k}=T_{i+k}$.
 Define a map $\gamma:[m+n]\times[m+n]\to\{-1,1\}$ by
-$\gamma(i,j)=\left\{\begin{array}[]{ll}1,&\mbox{if $i>j$},\\ -1,&\mbox{if $i\leq j$}.\end{array}\right.$
+$\gamma(i,j)=\left\{\begin{array}[]{ll}1&\mbox{if $i>j$},\\ -1&\mbox{if $i\leq j$}.\end{array}\right$.
 The following lemma will be used in the next section.
@@
 Proof.
-It can be verified by the braid relation $T_{1}T_{2}T_{1}=T_{2}T_{1}T_{2}$ case by case. ∎
+It can be verified directly by the braid relation $T_{1}T_{2}T_{1}=T_{2}T_{1}T_{2}$ case by case. ∎
 5.2. $q$-tensor superspaces
 We introduce a parity on $[m+n]$ by
-$\widehat{i}=\left\{\begin{array}[]{ll}0,&\mbox{if $1\leq i\leq m$},\\ 1,&\mbox{if $m<i\leq m+n$}\end{array}\right.$
+$\widehat{i}=\left\{\begin{array}[]{ll}0&\mbox{if $1\leq i\leq m$},\\ 1&\mbox{if $m<i\leq m+n$}\end{array}\right$.
 and let $q_{i}=q^{(-1)^{\widehat{i}}}$.
 Let $\mathbb{V}=\mathbb{V}_{\overline{0}}\oplus\mathbb{V}_{\overline{1}}$ be a vector space over $\mathbb{C}(q)$ with a specified basis $\{e_{1},e_{2},\ldots,e_{m+n}\}$, where $\mathbb{V}_{\overline{0}}$ (resp.
-$\mathbb{V}_{\overline{1}}$) is spanned by $\{e_{i}|1\leq i\leq m\}$ (resp.
-$\{e_{i}|m+1\leq i\leq m+n\}$).
+$\mathbb{V}_{\overline{1}}$) is spanned by $\{e_{i}~|~1\leq i\leq m\}$ (resp.
+$\{e_{i}~|~m+1\leq i\leq m+n\}$).
 For $I=(i_{d},\ldots,i_{2},i_{1})\in[m+n]^{d}$, we denote
@@
 Then $\{e_{I}~|~I\in[m+n]^{d}\}$ forms a basis of $\mathbb{V}^{\otimes d}$.
-As before, $\mathfrak{S}_{d}$ acts on $[m+n]^{d}$ from the right hand side.
+As before, $\mathfrak{S}_{d}$ acts on $m+n]^{d}$ from the right hand side.
 The right $\mathscr{H}_{d}$-action on $\mathbb{V}^{\otimes d}$ is defined on the basis element $e_{I}$ as follows: for $1\leq k<d$,
-(5.1) $e_{I}.T_{k}=\left\{\begin{array}[]{ll}(-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}},&\mbox{if $i_{k}>i_{k+1}$},\\ (-1)^{\widehat{i_{k}}}q_{i_{k}}e_{I},&\mbox{if $i_{k}=i_{k+1}$},\\ (-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}}+(q-q^{-1})e_{I},&\mbox{if $i_{k}<i_{k+1}$}.\end{array}\right.$
+(5.1) $e_{I}.T_{k}=\left\{\begin{array}[]{ll}(-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}}&\mbox{if $i_{k}>i_{k+1}$},\\ (-1)^{\widehat{i_{k}}}q_{i_{k}}e_{I}&\mbox{if $i_{k}=i_{k+1}$},\\ (-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}}+(q-q^{-1})e_{I}&\mbox{if $i_{k}<i_{k+1}$}.\end{array}\right$.
 We can rewrite (5.1) as
@@
 $\widetilde{T}(\mathbb{V})=\bigoplus_{d\geq 0}\widetilde{T}_{d}(\mathbb{V}).$
-Endow an algebra structure on $\widehat{T}(\mathbb{V})$ whose product $\cdot$ is defined by
+Endow an algebra structure on $\widetilde{T}(\mathbb{V})$ whose product $\cdot$ is defined by
 $(e_{j_{k}}\cdots e_{j_{1}}\otimes T_{\tau})\cdot(e_{i_{d}}\cdots e_{i_{1}}\otimes T_{\sigma})=e_{j_{k}}\cdots e_{j_{1}}e_{i_{d}}\cdots e_{i_{1}}\otimes T_{\tau}^{\uparrow d}T_{\sigma}.$
@@
 $L(\varphi):\widetilde{T}(\mathbb{V})$ $\longrightarrow$ $\widetilde{T}(\mathbb{V}),$
-[…]
+$\psi$~ $\longmapsto$ $\varphi\cdot\psi.$
 For $j\in[m+n]$, define $L(e_{j}^{*}):\widetilde{T}(\mathbb{V})\rightarrow\widetilde{T}(\mathbb{V})$ by
@@
 For $L(e_{j}^{*})$, we only need to show that it commutes with the action of $\mathscr{H}_{\infty}$.
-We introduce a linear map $h_{k}^{(j)}:T_{d}(\mathbb{V})\rightarrow\widetilde{T}_{d}(\mathbb{V})$ for $1\leq k\leq d$ defined by
+We introduce a linear map $h_{k}^{(j)}:\widetilde{T}_{d}(\mathbb{V})\rightarrow\widetilde{T}_{d}(\mathbb{V})$ for $1\leq k\leq d$ defined by
 $h_{k}^{(j)}(e_{i_{d}}\cdots e_{i_{1}})=(-1)^{\widehat{j}(\widehat{i_{d}}+\cdots+\widehat{i_{k+1}})}\langle e_{j}^{*},e_{i_{k}}\rangle g_{j}(e_{i_{d}})\cdots g_{j}(e_{i_{k+1}})f_{j}(e_{i_{k-1}})\cdots f_{j}(e_{i_{1}}).$
-It is just the $k$-th term of $L(e_{j}^{*})$ over $T_{d}(\mathbb{V})$ in (5.3).
+It is just the $k$-th term of $L(e_{j}^{*})$ on $\widetilde{T}_{d}(\mathbb{V})$ in (5.3).
 For any $T\in\mathscr{H}_{\infty}$, we define $\pi(T)\in\mathrm{End}_{\mathbb{C}}(\widetilde{T}(\mathbb{V}))$ by $\pi(T)(\varphi)=\varphi\cdot T$ for $\varphi\in\widetilde{T}(\mathbb{V})$.
 To prove the well-definedness of $L(e_{j}^{*})$, it suffices to show that $\pi(T_{1}),...,\pi(T_{d-1})$ commutes with $\sum_{k=1}^{d}h_{k}^{(j)}$.
-We have the following lemma:
 Lemma 5.2.
-Write $h_{k}$ for $h_{k}^{(j)}$. Then for $r=1,...,d-1$, we have
+Write $h_{k}$ for $h_{k}^{(j)}$. Then for r=1,,d1r=1,...,d-1, we have
   1. (i)
@@
 We start the induction by $\pi(T_{r})h_{r-1}=h_{r-1}\pi(T_{r})$ which is the result of the last paragraph.
-For $k<r-1$, we have
+For k<r1k<r-1, we have
 $\pi(T_{r})h_{k}$ $=\pi(T_{r})q_{j}^{2\gamma(j,i_{k+1})}\pi(T_{k}^{-\gamma(j,i_{k+1})})h_{k+1}\pi(T_{k}^{\gamma(i_{k},i_{k+1})})\delta_{j,i_{k}}$
@@
 $=\pi(T_{k}^{\gamma(i_{k},i_{k+1})})h_{k+1}+\pi(T_{k}^{\gamma(j,i_{k+1})})h_{k}-q_{j}^{-2\delta_{j,i_{k+1}}}\pi(T_{k}^{\gamma(j,i_{k+1})})h_{k}.$
-Adding the above two equations, we get the result. ∎
+Adding up the above two equations, we get the result. ∎
 5.5. General linear quantum supergroups
-The general linear quantum supergroup $\mathbf{U}_{m|n}=U_{q}(\mathfrak{gl}_{m|n})$ is a $\mathbb{C}(q)$-algebra generated by
-$\left\{\begin{array}[]{lll}\mbox{even generators:}&E_{i},F_{i},K_{j}^{\pm 1},&1\leq i,j\leq m+n,i\neq m,m+n;\\ \mbox{odd generators:}&E_{m},F_{m}&\end{array}\right.$
+The general linear quantum supergroup $U_{q}(\mathfrak{gl}(m|n))$ is a $\mathbb{C}(q)$-algebra generated by
+[…]
 subject to the following relations:
@@
 $F_{m-1,m+2}$ $=F_{m-1}F_{m}F_{m+1}-q^{-1}F_{m-1}F_{m+1}F_{m}-qF_{m}F_{m+1}F_{m-1}+F_{m+1}F_{m}F_{m-1}.$
-There is a Hopf superalgebra structure on $\mathbf{U}_{m|n}$, whose comultiplication $\Delta$ is
+There is a Hopf superalgebra structure on $U_{q}(\mathfrak{gl}(m|n))$, whose comultiplication $\Delta$ is
 $\Delta(K_{i})=K_{i}\otimes K_{i},$
@@
 $\Delta^{(d)}(F_{i})$ $=F_{i}\otimes 1^{\otimes d}+K_{i+1}K_{i}^{-1}\otimes\Delta^{(d-1)}(F_{i}).$
-Particularly, $|\Delta^{(d)}(E_{i})|=|E_{i}|$ and $|\Delta^{(d)}(F_{i})|=|F_{i}|$, where ||$\cdot$|| ⋅ | means the parity of elements in $\mathbf{U}_{m|n}$.
-The superspace $\mathbb{V}$ admits a natural representation of $\mathbf{U}_{m|n}$ via
+Particularly, $|\Delta^{(d)}(E_{i})|=|E_{i}|$ and $|\Delta^{(d)}(F_{i})|=|F_{i}|$, where $|\cdot|$ means the parity of elements in $U_{q}(\mathfrak{gl}(m|n))$.
+The superspace $\mathbb{V}$ admits a natural representation of $U_{q}(\mathfrak{gl}(m|n))$ via
 $K_{i}.e_{j}=q_{i}^{\delta_{ij}}e_{j},\quad E_{i}.e_{j}=\delta_{i+1,j}e_{i},\quad F_{i}.e_{j}=\delta_{i,j}e_{i+1}.$
-Then the $q$-tensor superspace $\mathbb{V}^{\otimes d}$ is also equipped with a $\mathbf{U}_{m|n}$-module structure via $\Delta^{(d)}$.
-We shall show that the $\mathbf{U}_{m|n}$-module structure can be extended over the $\widetilde{T}(\mathbb{V})$ as $x.(e_{I}\otimes T_{\sigma})=(x.e_{I})\otimes T_{\sigma}$ for $x\in\mathbf{U}_{m|n},e_{I}\in\mathbb{V}^{\otimes d},T_{\sigma}\in\mathscr{H}_{\infty}$.
+Then the $q$-tensor superspace $\mathbb{V}^{\otimes d}$ is also equipped with a $U_{q}(\mathfrak{gl}(m|n))$-module structure via $\Delta^{(d)}$.
+We shall show that the $U_{q}(\mathfrak{gl}(m|n))$-module structure can be extended to $\widetilde{T}(\mathbb{V})$ by $x.(e_{I}\otimes T_{\sigma})=(x.e_{I})\otimes T_{\sigma}$ for any $x\in U_{q}(\mathfrak{gl}(m|n)),e_{I}\in\mathbb{V}^{\otimes d},T_{\sigma}\in\mathscr{H}_{\infty}$.
 It can be seen instantly for $K$, since
@@
 For any $i,j\in[m+n]$ and $1\leq k<d$, as operators on $\widetilde{T}(\mathbb{V})$, we have
-$K_{j}L(e_{i})=q_{j}^{\delta_{ij}}L(e_{i})K_{j},\quad K_{j}L(e_{i}^{*})=q_{j}^{-\delta_{ij}}L(e_{i}^{*})K_{j},\quad K_{j}L(T_{k})=L(T_{k})K_{j}$
+$K_{j}L(e_{i})=q_{j}^{\delta_{ij}}L(e_{i})K_{j},\quad K_{j}L(e_{i}^{*})=q_{j}^{-\delta_{ij}}L(e_{i}^{*})K_{j},\quad K_{j}L(T_{k})=L(T_{k})K_{j}.$
 Proof.
@@
 Now let us prove (6.2), which is equivalent to
-[…]
+(6.4) $L(e_{i}^{*})L(e_{j}^{*})|_{\widetilde{T}_{d}(\mathbb{V})}=(-1)^{\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(T_{1}^{\gamma(i,j)})L(e_{j}^{*})L(e_{i}^{*})|_{\widetilde{T}_{d}(\mathbb{V})},$
 for $d=0,1,2,\ldots$.
@@
 If $i_{d}=j$, we can get the result similarly. ∎
-6.2. The representations of $U_{m|n}$ on $\widehat{T}(\mathbb{V})$
+6.2. The representations of $U_{q}(\mathfrak{gl}(m|n))$ on $\widetilde{T}(\mathbb{V})$
 For $1\leq i<m+n$ and $d\in\mathbb{N}$, we denote
@@
 Proof.
-For $I=(i_{d-1},\ldots,i_{2},i_{1})\in[m+n]^{d-1},k\in[m+n]$, by direct calculation, we have
+For $I=(i_{d-1},\ldots,i_{2},i_{1})\in[m+n]^{d-1}$ and $k\in[m+n]$, by direct calculation, we have
 $A_{i}^{(d)}(e_{k}e_{i_{d-1}}\cdots e_{i_{1}})$
@@
 Proof.
-We only prove the first equation. The second one can be proved similarly. We use induction on $d$. It is clear for $d=1$. Suppose the equation holds for $d=r-1$, then
+We only prove the first equation. The second one can be proved similarly. We use induction on $d$. It is clear for $d=1$. Suppose the equation holds for d=r1d=r-1, then
 $E_{i}K_{i}^{-1}(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
@@
 $=(A_{i}^{(1)}\otimes K_{i+1}^{-1}+(-1)^{\delta_{im}\widehat{i_{r}}}K_{i}^{-1}\otimes A_{i}^{(r-1)})(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
-$=A_{i}^{(r)}(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})=L(e_{i})L(e_{i+1}^{*})(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}}),$
-where the forth equality follows from the inductive hypothesis, and the sign comes from the super-action of $E_{i}K_{i}^{-1}$.
-We further denote $E_{i,i+1}=E_{i}$, $E_{i+1,i}=F_{i}$ and
+$=A_{i}^{(r)}(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
+$=L(e_{i})L(e_{i+1}^{*})(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}}),$
+where the forth equality follows from the inductive hypothesis, and the sign comes from the super-action of $E_{i}K_{i}^{-1}$.∎
+We further denote $E_{i,i+1}=E_{i}$, $E_{i+1,i}=F_{i}$, and
 $E_{ij}=E_{ik}E_{kj}-q_{k}E_{kj}E_{ik},\quad E_{ji}=E_{jk}E_{ki}-q_{k}^{-1}E_{ki}E_{jk},\quad(i<k<j).$
@@
 As above, we only prove the first equation.
-We use induction on $j-i$.
-It has been showed in Lemma 6.5 for the case $j-i=1$.
-On $\mathbb{V}^{\otimes d}$, suppose $E_{ik}K_{i}^{-1}=L(e_{i})L(e_{k}^{*}),i<k<j$.
+We use induction on jij-i.
+It has been showed in Lemma 6.5 for the case ji=1j-i=1.
+On $\mathbb{V}^{\otimes d}$, suppose $E_{ik}K_{i}^{-1}=L(e_{i})L(e_{k}^{*})$ for $i<k<j$.
 Then
@@
 Remark 6.7.
-Based on these identities on $\mathbb{V}^{\otimes d}$ and the fact that $L(e),L(e^{*}),K,K^{-1}$ all commute with $\mathscr{H}_{\infty}$, we can actually have a well-defined $\mathbf{U}_{m|n}$-module structure on $\widetilde{T}(\mathbb{V})$ and get the same result.
+Based on these identities on $\mathbb{V}^{\otimes d}$ and the fact that $L(e),L(e^{*}),K,K^{-1}$ all commute with $\mathscr{H}_{\infty}$, we can actually have a well-defined $U_{q}(\mathfrak{gl}(m|n))$-module structure on $\widetilde{T}(\mathbb{V})$ and get the same result.
 Furthermore, we can get the following proposition.
@@
 For any $v_{1},...,v_{k}\in\mathbb{V}$, $v_{1}^{*},...,v_{k}^{*}\in\mathbb{V}^{*}$, we have
-$L(v_{k})\cdots L(v_{1})L(v_{1}^{*})\cdots L(v_{k}^{*})\in\rho(\mathbf{U}_{m|n}),$
-where $\rho$ denotes the representation of $U_{m|n}$ on $\widetilde{T}(\mathbb{V})$.
+$L(v_{k})\cdots L(v_{1})L(v_{1}^{*})\cdots L(v_{k}^{*})\in\rho(U_{q}(\mathfrak{gl}(m|n))),$
+where $\rho$ denotes the representation of $U_{q}(\mathfrak{gl}(m|n))$ on $\widetilde{T}(\mathbb{V})$.
 Proof.
@@
 $L(e_{i})L(e_{j})L(e_{k}^{*})=aL(e_{i})L(e_{k}^{*})L(e_{j})+bL(e_{j})L(e_{k}^{*})L(e_{i})+cK_{k}^{-1}L(e_{i})+dK_{k}^{-1}L(e_{j}),$
-where $a,b,c,d\in\mathbb{C}(q)$. Recall from Proposition 6.1 and 6.3 that
+where $a,b,c,d\in\mathbb{C}(q)$. Recall from Propositions 6.1 and 6.3 that
 $L(e_{i})L(e_{j})L(T^{-\gamma(k,j)})L(e_{k}^{*})$ $=(-1)^{\widehat{j}\widehat{k}}(L(e_{i})L(e_{k}^{*})L(e_{j})-\delta_{jk}L(e_{i})K_{j}^{-1})$
@@
 $=(-1)^{\widehat{i}\widehat{j}}L(e_{j})L(e_{i})L(T^{-\gamma(k,i)})L(e_{k}^{*})+c_{1}L(e_{i})L(e_{j})L(T^{-\gamma(k,j)})L(e_{k}^{*})+c_{1}c_{2}L(e_{i})L(e_{j})L(e_{k}^{*}),$
-where $c_{1},c_{2}$ is ±$\pm(q-q^{-1})$ or 00.
+where $c_{1},c_{2}$ is $\pm(q-q^{-1})$ or 0.
 By direct calculation, we have $c_{1}=0$ or $c_{2}=0$ (i.e., $c_{1}c_{2}=0$), since $\gamma(i,j)=-\gamma(k,i)$ or $\gamma(j,i)=-\gamma(k,j)$.
@@
 For any nonnegative integer $k$, we denote by
-$[k]=\frac{q^{k}-q^{-k}}{q-q^{-1}}\quad\mbox{and}\quad[k]!=[1][2]\cdots[k]$
+$k]=\frac{q^{k}-q^{-k}}{q-q^{-1}}\quad\mbox{and}\quad[k]!=[1][2]\cdots[k]$
 the $q$-quantum integers and $q$-quantum factorials, respectively.
 For $i\in[m+n],a\in\mathbb{Z}$, set
-$[K_{i}:a]=\frac{q_{i}^{a}K_{i}-q_{i}^{-a}K_{i}^{-1}}{q_{i}-q_{i}^{-1}},$
-$[K_{i}]^{h}_{!}=[K_{i}:0][K_{i}:-1]\cdots[K_{i}:1-h],\quad\mathrm{for}\ h\in\mathbb{N_{+}},$
-and $[K_{i}]^{0}_{!}=1$. Then by Lemma 6.2 we have
+$K_{i}:a]=\frac{q_{i}^{a}K_{i}-q_{i}^{-a}K_{i}^{-1}}{q_{i}-q_{i}^{-1}},$
+$K_{i}]^{h}_{!}=[K_{i}:0][K_{i}:-1]\cdots[K_{i}:1-h],\quad\mathrm{for}\ h\in\mathbb{N_{+}},$
+and $K_{i}]^{0}_{!}=1$. Then by Lemma 6.2 we have
 (7.1) $\underbrace{L(e_{i})\cdots L(e_{i})}_{h}\underbrace{L(e_{i}^{*})\cdots L(e_{i}^{*})}_{h}=[K_{i}]^{h}_{!}.$
-Set =$\mathcal{I}=\{(i_{d},i_{d-1},...,i_{1})\in[m+n]^{d}~{}|~{}i_{d}\leq i_{d-1}\leq\cdots\leq i_{1}\}$.
-For any $I\in\mathcal{I}$, let $[\mathbf{m}(I)]!=[\mathbf{m}_{1}(I)]![\mathbf{m}_{2}(I)]!\cdots[\mathbf{m}_{m+n}(I)]!$ where $\mathbf{m}_{k}(I)$ is the multiplicity of $k$ in $I$.
+Set $\mathcal{I}=\{(i_{d},i_{d-1},...,i_{1})\in[m+n]^{d}~|~i_{d}\leq i_{d-1}\leq\cdots\leq i_{1}\}$.
+For any $I\in\mathcal{I}$, let $\mathbf{m}(I)]!=[\mathbf{m}_{1}(I)]![\mathbf{m}_{2}(I)]!\cdots[\mathbf{m}_{m+n}(I)]!$ where $\mathbf{m}_{k}(I)$ is the multiplicity of $k$ in $I$.
 By (7.1) and Proposition 6.1 we have
@@
 Proof.
-Note that, for any $I,J\in\mathcal{I}$, we have
-$[K_{i}]^{\mathbf{m}_{i}(I)}_{!}.e_{J}=\begin{cases}[\mathbf{m}_{i}(I)]!e_{J},&\text{if }\mathbf{m}_{i}(I)=\mathbf{m}_{i}(J);\\[3.0pt] 0,&\text{if }\mathbf{m}_{i}(I)>\mathbf{m}_{i}(J).\end{cases}$
-Then by (7.2), one has
+For any $I,J\in\mathcal{I}$, if $\mathbf{m}_{i}(I)=\mathbf{m}_{i}(J)$, we have
+$K_{i}]^{\mathbf{m}_{i}(I)}_{!}.e_{J}=[\mathbf{m}_{i}(I)]!e_{J},$
+and if $\mathbf{m}_{i}(I)>\mathbf{m}_{i}(J)$, we have
+$K_{i}]^{\mathbf{m}_{i}(I)}_{!}.e_{J}=0.$
+In the case where $\mathbf{m}_{i}(I)<\mathbf{m}_{i}(J)$ for some $i$, there must exist some index $j$ such that $\mathbf{m}_{j}(I)>\mathbf{m}_{j}(J)$. Thus by (7.2), one has
 $L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{*})\cdots L(e_{i_{d}})(e_{J})=\delta_{I,J}\Bigg(\prod_{i\in[m+n]}[\mathbf{m}_{i}(I)]!\Bigg)e_{J}=\delta_{I,J}[\mathbf{m}(I)]!e_{J}.$
@@
 7.2. Descriptions of commutants
-For any $d\geq 1$ and $0\leq k\leq d$, we denote by $\mathcal{A}_{d}^{k}$ the subspace of $\mathrm{End}_{\mathbb{C}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ spanned by operators of the form
+For any $d\geq 1$ and $0\leq k\leq d$, we denote by $\mathcal{A}_{d}^{k}$ the subspace of $\mathrm{End}_{\mathbb{C}(q)}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ spanned by operators of the form
 $L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{*})\cdots L(v_{d}^{*}),$
 with $w_{1},\ldots,w_{d}\in\mathbb{V}$, $\sigma\in\mathscr{H}_{d+k}$, and $v_{1}^{*},\ldots,v_{k}^{*}\in\mathbb{V}^{*}$.
-For any $T\in\mathscr{H}_{d+k}$, we define $\pi_{d+k}(T)\in\mathrm{End}_{\mathbb{C}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ by $\pi_{d+k}(T)(\varphi)=\varphi\cdot T$ where $\varphi\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$.
-Denote $\mathcal{B}_{d+k}=\pi_{d+k}(\mathscr{H}_{d+k})$.
+The representation $\pi_{d+k}$ of $\mathscr{H}_{d+k}$ on $\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$ is given by right multiplication and we identify $\pi_{d+k}(\mathscr{H}_{d+k})$ with $\mathscr{H}_{d+k}$.
 Theorem 7.2.
-Asumme that $[d]!\neq 0$. We have
-$\mathrm{Comm}(\mathcal{A}_{d}^{k})=\mathcal{B}_{d+k},$
+Asumme that $d]!\neq 0$. We have
+$\mathcal{A}_{d}^{k}=\mathrm{End}_{\mathscr{H}_{d+k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}),$
 and
-$\mathrm{Comm}(\mathcal{B}_{d+k})=\mathcal{A}_{d}^{k}.$
+$\mathrm{End}_{\mathcal{A}_{d}^{k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})=\mathscr{H}_{d+k}.$
 Proof.
-It is easy to see that $\mathcal{A}_{d}^{k}\subset\mathrm{Comm}(\mathcal{B}_{d+k})$.
-Since $\mathcal{H}_{d}(q)$ is semisimple, by the dobule commutant theorem, it suffices to show $\mathrm{Comm}(\mathcal{B}_{d+k})\subset\mathcal{A}_{d}^{k}$.
-For $f\in\mathrm{Comm}(\mathcal{B}_{d+k}),\phi\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have
+It is easy to see that $\mathcal{A}_{d}^{k}\subset\mathrm{End}_{\mathscr{H}_{d+k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$.
+Since $\mathcal{H}_{d}(q)$ is semisimple, by the dobule commutant theorem, it suffices to show $\mathrm{End}_{\mathscr{H}_{d+k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})\subset\mathcal{A}_{d}^{k}$.
+For $f\in\mathrm{End}_{\mathscr{H}_{d+k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}),\phi\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have
 $f(\phi)$ $=f(A_{d}(\phi))$
@@
 $=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}L(f(e_{i_{d}}\cdots e_{i_{1}}))L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})\phi,$
-where $\tau_{I}=L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})\phi\in\mathscr{H}_{d+k}$.
-So, $f=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}L(f(e_{i_{d}}\cdots e_{i_{1}}))L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})$.
-Since $L(f(e_{i_{d}}\cdots e_{i_{1}}))\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have $f\in\mathcal{A}_{d}^{k}$.
-∎
-Theorem 7.3 (Mitsuhashi).
-Asumme that $[d]!\neq 0$. For each $d\geq 1$, we have
-$\mathrm{Comm}(\pi_{d}(\mathscr{H}_{d}(q)))=\rho_{d}(\mathbf{U}_{m|n}),$
+where $\tau_{I}=L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})\phi\in\mathscr{H}_{d+k}$. So
+$f=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}L(f(e_{i_{d}}\cdots e_{i_{1}}))L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*}).$
+Since $L(f(e_{i_{d}}\cdots e_{i_{1}}))\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have $f\in\mathcal{A}_{d}^{k}$. ∎
+Corollary 7.3 (Mitsuhashi).
+Asumme that $d]!\neq 0$. For each $d\geq 1$, we have
+$\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))=\mathrm{End}_{\mathscr{H}_{d}}(\mathbb{V}^{\otimes d}),$
 and
-$\mathrm{Comm}(\pi_{d}(\mathscr{H}_{d}(q)))=\rho_{d}(\mathbf{U}_{m|n}),$
-where $\rho_{d}$ denotes the representation of $\mathbf{U}_{m|n}$ on $\mathbb{V}^{\otimes d}\cong\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d}$.
+$\mathrm{End}_{\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))}(\mathbb{V}^{\otimes d})=\mathscr{H}_{d},$
+where $\rho_{d}$ denotes the representation of $U_{q}(\mathfrak{gl}(m|n))$ on $\mathbb{V}^{\otimes d}\cong\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d}$.
 Proof.
-We just need to prove that $\rho_{d}(\mathbf{U}_{m|n})=\mathcal{A}_{d}^{0}$.
-Proposition 6.8 shows that $\mathcal{A}_{d}^{0}\subset\rho_{d}(\mathbf{U}_{m|n})$.
+We just need to prove that $\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))=\mathcal{A}_{d}^{0}$.
+Proposition 6.8 shows that $\mathcal{A}_{d}^{0}\subset\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))$.
 The other side is proved as follows.
-We first notice that $K_{i}\in\mathrm{Comm}(\pi_{d}(\mathscr{H}_{d}(q)))$.
-In the proof of Theorem 7.2, we have seent that $K_{i}\in\mathcal{A}_{d}^{0}$.
+We first notice that $K_{i}\in\mathrm{End}_{\mathscr{H}_{d}}(\mathbb{V}^{\otimes d})$.
+In the proof of Theorem 7.2, we have seen that $K_{i}\in\mathcal{A}_{d}^{0}$.
 Actually, we have
 $K_{i}=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}q_{j}^{\delta_{ii_{d}}+\cdots+\delta_{ii_{1}}}L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*}).$
-Since Theorem 6.6 shows that $E_{ij}=L(e_{i})L(e_{j}^{*})K_{i}$ and $E_{ij}=K_{j}^{-1}L(e_{j})L(e_{i}^{*})$ for $i<j$ on ${\mathbb{V}^{\otimes d}}$, these elements commute with $\mathcal{B}_{d}$, then $\rho_{d}(\mathbf{U}_{m|n})\subset\mathcal{A}_{d}^{0}$.
+It follows from Theorem 6.6 that $E_{ij}=L(e_{i})L(e_{j}^{*})K_{i}$ and $E_{ij}=K_{j}^{-1}L(e_{j})L(e_{i}^{*})$ for $i<j$ on $\mathbb{V}^{\otimes d}$}.
+So these elements commute with $\mathcal{B}_{d}$, and hence $\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))\subset\mathcal{A}_{d}^{0}$.
 ∎